### v1

# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

Abstract
The diffusion probability model is a state-of-the-art generative model that generates an image by applying a neural network iteratively. Moreover, this generation process is regarded as an algorithm solving a diffusion ordinary differential equation (ODE) or stochastic differential equation (SDE). Based on the analysis of the truncation error of the diffusion ODE and SDE, our study proposes a training-free algorithm that generates high-quality 512 x 512 and 1024 x 1024 images in eight steps, with flexible guidance scales. To the best of my knowledge, our algorithm is the first one that samples a 1024 x 1024 resolution image in 8 steps with an FID performance comparable to that of the latest distillation model, but without additional training. Meanwhile, our algorithm can also generate a 512 x 512 image in 8 steps, and its FID performance is better than the inference result using state-of-the-art ODE solver DPM++ 2m in 20 steps. The result of our algorithm in generating high-quality 512 x 512 images and 1024 x 1024 images with five-step and six-step inference is also comparable to the latest distillation model. Moreover, unlike most distillation algorithms, which achieve state-of-the-art FID performance by fixing the sampling guidance scale, and which sometimes cannot improve their performance by adding inference steps, our algorithm uses a flexible guidance scale on classifier-free guidance sampling. The increase in inference steps enhances its FID performance. Additionally, the algorithm can be considered a plug-in component compatible with most ODE solvers and latent diffusion models. Extensive experiments are performed using the COCO 2014, COCO 2017, and the LAION dataset. Specifically, we validate our eight-step image generation algorithm using the COCO 2014, COCO 2017, and LAION validation datasets with a 5.5 guidance scale and a 7.5 guidance scale, respectively. Furthermore, the FID performance of the image synthesis in 512 x 512 resolution with a 5.5 guidance scale is 15.7, 22.35, and 17.52, meaning it is comparable with the state-of-the-art ODE solver DPM++ in 20 steps, whose best FID performance is 17.3, 23.75, and 17.33, respectively. Further, it also outperforms the state-of-the-art AMED-plugin solver, whose FID performance is 19.07, 25.50, and 18.06. We also apply the algorithm in five-step inference without additional training, for which the best FID performance of our algorithm in COCO 2014, COCO 2017, and LAION is 19.18, 23.24, and 19.61, respectively, which is comparable to the performance of the state-of-the-art AMED Pulgin solver in eight steps, SDXL-turbo in four steps, and the state-of-the-art diffusion distillation model Flash Diffusion in five steps. Then, we validate our algorithm in synthesizing 1024 * 1024 images, whose FID performance in COCO 2014, COCO 2017, and LAION using eight-step inference is 17.84, 24.42, and 19.25, respectively. Thus, it outperforms the SDXL-lightning in eight steps, Flash DiffusionXL in eight steps, and DMD2 in four steps. Moreover, the FID performance of the six-step inference of our algorithm in the 1024 x 1024 image synthesis is 23, which only has a limited distance to the state-of-the-art distillation model mentioned above. We also use information theory to explain the advantage of our algorithm and why it achieves a strong FID performance.
1 Introduction
The diffusion probability model (DM)[12], a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision. A typical diffusion model, such as stable diffusion (SD)[28], is inspired by the Brownian diffusion process in physics. It consists of two different processes: diffusion and reverse. The diffusion process adds small variants of Gaussian noise to the image step by step to facilitate its gradual conversion into pure Gaussian noise, while the reverse process uses the score matching[37] algorithm to train a neural network studying the gradient of the diffusion process by denoising noisy images[25]. Then, the reverse of the diffusion process occurs. The diffusion process can be described by a diffusion stochastic differential equation (SDE) whose form is similar to the Brownian diffusion SDE. Moreover, the reverse process can be described by an ordinary differential equation (ODE) or an SDE[37]. The ODE is equivalent to the SDE in the reverse process, which generates random variables step by step by transporting the standard Gaussian distribution to a target probability distribution by changing its geometric shape. However, it requires fewer inference steps to sample a high-quality image than solving the SDE, as it prevents involving a random effect during sampling. However, solving diffusion ODEs still requires a large amount of computational resources. Hence, solving this problem remains an active research area. Research related to it can be categorized into two classes. The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process[12, 24, 35, 45, 15, 21]. The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and it reduces the inference steps by training a small distillation model[26, 36, 19, 44, 43, 31, 32, 16, 38, 22]. The training-free method typically necessitates a 20-step inference[24], while algorithms requiring additional training can generate an image within four to eight steps. However, our results show that by selecting the proper hyperparameter of the ODE solver and adding a training free diffusion model decorator to exploit the capability of the latent diffusion model, the inference speed can exceed expectations. To that end, we propose a new inference method that can generate 512 x 512 images in eight steps without additional training, and its Frechet Inception Distance (FID)[10] performance surpasses that of the state-of-the-art ODE solver in 20 steps. Furthermore, when applying our algorithm in the five-step inference, it also outperforms the state-of-the-art distillation algorithm in the same number of steps. We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step inference exceeds most of the latest distillation models in eight steps. We perform extensive experiments using the COCO 2014[20], COCO 2017 [20], and LAION datasets[33], and exhibit the generation result in Fig 1. We also use information theory to explain why our algorithm achieves a strong FID result.
Figure 1: The comparison between our algorithm and dpm++ solver in few step inference
2 Related Work
2.1 Diffusion ODE
The image generation process that uses a diffusion model can be regarded as solving a special diffusion SDE or ODE. A typical diffusion model has two processes: forward and reverse. The forward process perturbs the image by adding Gaussian noise to it, which can be described by the following SDE:
$dx=f(t)xdt+g(t)dw,$ (1)
where the $w$ is a standard Wiener process. Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms. In this study, we use the variant preserve (VP) SDE[37] to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm[13]. Specifically, its choice is as follows:
$f(t)=-\frac{1}{2}\beta(t),$ (2)
$g(t)=\sqrt{\beta(t)},$ (3)
$\beta(t)=\beta_{min}+t(\beta_{max}-\beta_{min}).$ (4)
By solving the Eq. 1 step by step with a given $x_{0}$, we obtain a noisy image $x_{t}$. Finally, when applying iteration in $T$ step, we obtain $x_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}I)$. Additionally, using the reparameterization algorithm, we can sample $x_{t}$ within one step[12].
In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE[15]:
$\frac{dx}{dt}=\frac{\dot{s}(t)}{s(t)}x_{t}+s(t)^{2}\dot{\sigma}(t)\epsilon_{\theta}(x_{t},t,c),$ (5)
where $c$ is the condition of input, like the caption that describes the image, $\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)s_{\theta}(x_{t},c)$, where $s_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)}$. Moreover, the score function normally used in VP SDE, it is trained by Eq. 6
[…]
and $q_{t}$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$.
Specifically, the $s(t)$ and $\sigma(t)$ is defined by Eq. 7 and Eq. 8:
$\sigma(t)=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1},$ (7)
$s(t)=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}},$ (8)
where $\beta_{d}=\beta_{max}-\beta_{min}$.
Using Eq. 5 as a reverse ODE can help exploit the advantage of adapting the DDPM algorithm to the common diffusion framework proposed by Karras [15]. However, the algorithm implementation should also consider whether a neural network can study this input–output pair if the input and output in different timesteps have an obvious gap in the variance when using this framework. So, to let the neural network study properly, we should formulate the equation as
$\epsilon_{\theta}(x_{t},t,c)=c_{out}(\sigma_{t})F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$ (9)
where $F_{\theta}$ is the real function that the neural network studies, $c_{in}$ scales the input to reduce its variance, $c_{out}$ scales the output to use the common framework proposed by Karras, and $c_{noise}(\sigma)$ transforms the variance $\sigma$ to the condition input $t$ of the neural network.
For simplicity, the following section will use $D_{\theta}$ to denote the centrifuge of the possible denoised images predicted by the neural network when given a noisy image $x_{t}$:
$D_{\theta}(x_{t};\sigma)=$ $c_{skip}(\sigma)x_{t}+$ (10)
$c_{out}(\sigma)F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$
where $c_{skip}(\sigma)=1$ in DDPM.
2.2 Diffusion ODE Solver
Sampling an image from a diffusion model is equivalent to solving the Eq. 5. Moreover, different solvers have been considered in previous research[15, 23, 24]. The inference process approximates the trajectory of the reverse ODE by discretizing its path. The research on reverse ODE solvers attempts to precisely estimate the next step, thereby minimizing the truncation error caused by the discretization. As the culmination of the truncation error will affect the result of synthesis, minimizing the truncation error in each step is equivalent to subtracting the demand for an increased inference step that has caused precise discretization.
Among those solvers, the DPM++ solver[24] is the latest state-of-the-art solver designed specifically for classifier-free guidance (CFG)[11] sampling. It predicts the next step via the following formula:
$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}D_{\theta}(x_{t(\lambda)};\sigma_{t(\lambda)})d\lambda,$ (11)
where $\lambda:=\log{a_{t}}-\log{\sigma_{t}}$ is a log signal-to-noise ratio (SNR) of the noisy image, while the function $(\lambda)$ transforms $\lambda$ to the corresponding $t$.
We use the Taylor formula to obtain an approximation of the Eq. 11, and then there is a first-order DPM++1s solver:
$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}e^{h}D_{\theta}(x_{s},\sigma_{s}),$ (12)
in which $h=\lambda_{t}-\lambda_{s}$.
Moreover, by using a multistep method, we obtain the second-order DPM++2m solver:
$x_{t_{i}}=\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}x_{t_{i-1}}-\alpha_{t_{i}}(e^{-h_{i}}-1)D_{i},$ (13)
where $D_{i}$ is as follows:
$D_{i}$ $=(1+\frac{1}{2r_{i}})D_{\theta}(x_{t_{i-1}};\sigma_{t_{i-1}})$ (14)
$-\frac{1}{2r_{i}}D_{\theta}(x_{t_{i-2}};\sigma_{t_{i-2}}),$
and $r_{i}=\frac{h_{i-1}}{h_{i}}$.
2.3 Diffusion Distillation
The training-free method, which optimizes the truncation error in each step to minimize the demand for precise discretion. Unlike this method, the distillation algorithm[19, 32, 26, 36, 31] considers the whole inference process of applying a neural network iteratively as a large neural network. The distillation algorithm, which normally trains a small neural network, entails reducing its inference steps. However, training a neural network with high-quality content, a fast inference speed, and no additional parameters is difficult. Thus, the distillation algorithm, through the analysis of information theory[17] and experiments, can sometimes reduce the diversity of the generated images, thus reducing the FID performance. This approach may pose a potential risk in the future. AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules. Moreover, when people draw an image via their craft, they will sometimes be affected by the image they have previously watched. If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such generation will not be deemed plagiarism. However, generative models that use Internet images as training data are always denied by people who claim that if those models generate similar content in their dataset, it should be considered plagiarism. Furthermore, diversity is among the metrics that affect people’s impression of generative content.
2.4 Latent Diffusion Model
Another method for reducing the calculation resources of the inference is to encode images into a low-dimensional latent space. Stable diffusion[28] performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image to use the attention mechanism in the U-Net[29] efficiently. One team’s experiment[28] shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism[40]. In contrast, its perceptual information, which makes the image appear crisp, is strongly linked to local regions and can be modeled by a convolutional neural network (CNN)[18]. Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder[2] to let the following inference ignore the perceptual detail. Then, it performs diffusion inference in latent space via a U-Net[29] with an attention mechanism to model the semantic information. Experiments show that compared with vector quantify VAE[41], the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.
3 Approach
3.1 Algorithm Based on Truncation Error Analysis
The basic logic of the diffusion ODE is as follows: When the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as detrimental as the non-noisy steps. These insights have been proved in a previous paper[15] via an experiment that discretizes the inference process by choosing a uniform variant gap in each discretized time step to obtain arrays: $\{\sigma_{i}=\frac{i}{N-1}\sigma_{max};i<64\}$ and $\{t_{i};t(\sigma_{i})\}$. Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$. The experimental results show that the truncation error increases monotonically as the start variant decreases.
[…]
Based on this observation, one researcher [15] proposes a following discrete method:
$\sigma_{i}=(\sigma_{min}^{\frac{1}{p}}+\frac{i}{N-1}(\sigma_{min}^{\frac{1}{p}}-\sigma_{max}^{\frac{1}{p}}))^{p}.$ (15)
Eq. 15 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part. Moreover, according to the observation from the paper[17], if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget. Considering the U-Net has 860 million parameters, and the variant of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical. Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
Based on that observation, one study [17] proposes another discrete method:
$t_{i}=(t_{min}^{\frac{1}{p}}+\frac{i}{N-1}(t_{min}^{\frac{1}{p}}-t_{max}^{\frac{1}{p}}))^{p}.$ (16)
When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part is larger than that in Karras’s original method.
Furthermore, we should consider the meaning of the score function by rewriting the Eq. 6 as follows:
$\mathcal{L}_{\theta}=\frac{E_{x_{0}}[q(x_{t}|x_{0})||s_{\theta}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}]}{E_{x_{0}}[q(x_{t}|x_{0})]},$ (17)
which means the score function is pointing to the average of all possible images that can cause this noisy image. Naturally, one can surmise that $\nabla_{x}\log q(x_{t})\approx\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage. In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
Based on the aforementioned idea and the experiment results in Section 4, we propose the following discrete method that aims to correct the output in the less noisy stage:
$t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}},$ (18)
where $\sigma_{stop}$ is defined via the following formula:
$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{N+stop+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (19)
which is not zero, as the decode $\beta$-VAE trained with a small KL-term constraint can handle a small amount of noise, and this setting can reduce the one-step inference budget. The visual different between our discrete method and the method proposed by the Karras’ paper is in Fig. 2
Figure 2: The comparison between our time schedule and the time schedule proposed by Karras
Additionally, consider the Eq. 17; the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, which can be achieved by modifying the skip connection and the backbone feature of the U-Net. Previous study[34] adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability. Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the special part of the noisy backbone feature map that lower than a particular frequency. The decorator only applies to the first two skip connections, and the scaler affected the backbone feature is $b_{1}$ and $b_{2}$, while the scaler affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
Also, consider that the normal ResNet[9] can be viewed as an ODE[5] by using the following formula:
$h_{t+1}=h_{t}+f_{\theta}(h_{t},t),$ (20)
$\frac{dh}{dt}=F_{\theta}(h_{t},t),$ (21)
where $h_{t}$ is the feature output in layer $t$.
The U-Net, as a black box ODE solver, can be regarded as another ODE. Thus, modifying the skip connection and scale backbone feature will further increase its output and the moving distance of the diffusion reverse ODE, which let the start step of using U-Net decorator should be parameterized in 1024 x 1024 image synthesis, as it contributes to the increase of moving distance in the noisy part. This behavior causes the global truncation error in solving ODE, which may let the variables be out of the distribution (OOD). We also visualize the decoded feature in the trajectory of different samplers in Fig. 3.
Figure 3: The comparison between our sampler and the original DPM++1s sampler. The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler. The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
We then visualize the decode trajectory of using different discrete methods by applying the same sampler DPM++ 1s sampler in Fig. 5. The normal time schedule using Free-U will generate a trajectory similar to the method using Karras’ time schedule without using Free-U, which exhibits the evidence of our previous claim.
Figure 4: The comparison between different discrete method when using the same DPM++ 1s Sampler in a few-step inference. The first line exhibit the trajectory of the normal Karras discrete method. The second one exhibit the trajectory of the Karras discrete method + Free-U. The third exhibite the normal discrete method with normal U-Net, and the final one exhibit the normal discrete method with Free-U. And the final line is similiar to the first line.
Our algorithm can be described by the Algorithm 1
Data: number of inference step $N$, initiated noise $\epsilon$, stop step in karras schedule $t_{stop}$, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$
Result: final result image
1 $\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$;
2 $ts=\{t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}};0\leq i<N\}$ $x=\epsilon$ while $t$ in tsts do
3 if $t$ equal to $t_{aug}$ then
4    $freeU(F_{\theta})$;
5
6   end if
7 $x=F_{\theta}(x)$
8 end while
Algorithm 1 custom Karras scheduler in the common framework of Diffusion ODE solving
3.2 Information Theory Analysis of the Different Algorithms
A common point of confusion in deep learning at an early age is how neural networks with so many parameters still have generalization ability[1]. The normal statistical model uses an extremely limited number of parameters to ensure that it can concisely describe the phenomenon and capture the essence of the observed data. For this problem, the study theory[8] claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions. Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{n}$. According to the information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can gain access to is $2^{H(D)}$[42]. Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$. The information bottleneck theory[39] shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ project to the $Z$. Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses. As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) [7], which face mode collapse[6] that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint [14]. One study[17] shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information. The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE)[2], in which, when $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bounce of the mutual information constraint will be abated, and there will be an increase in the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$. From this perspective, the diffusion model can be deemed as a hierarchy VAE[25], while progressive distillation[31] increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$. Where $X_{\theta}$ is a noisy feature output by solving ODE. Hence, the PSNR also increases. A recent example of diffusion distillation is SDXL-lightning[19], which adds GAN loss to the progressive distillation training, and its FID remains similar to normal progressive distillation.
However, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will also affect the mutual information, even if the output does not have an increased PSNR. We provide further analysis regarding this.
The diffusion ODE can be regarded as a Markov process $X_{\theta}(T)-X_{\theta}(T-1)-...-X_{\theta}(t)-...X_{\theta}(0)$, in which $X_{\theta}(t)$ follows the input of the diffusion model predicted by the U-Net, and we have the following Eq. 23:
$I(X_{\theta}(0),X_{\theta}(t))=H(X_{\theta}(0))-H(X_{\theta}(0)|X_{\theta}(t))$ (22)
$\leq H(X(0))-E_{X_{\theta}(t)}[p(X_{\theta}(0)|X_{\theta}(t),$
$*\log p_{\theta}(X_{\theta}(0)|X_{\theta}(t)))],$ (23)
in which $H(X(0))$ is the real $X(0)$ sample from the dataset.
In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$ while enhancing the log likelihood of the final synthesis. Moreover, Eq. 23 shows that if we modify the $X_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase. Most distillation methods that do not increase the PSNR of the output will still face this problem. We evaluate the Precision and Recall of Distribution (PRD)[30] of flash diffusion and our sampler to further validate this claim.
4 Experiments
4.1 Performance
In this study, we employ $p_{1}=7$, $p_{2}=1.2$, $stop=3$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images. For the eight-step inference, as the precise discretization method, we can completely avoid using free-U, which can enhance the aesthetic of images while reducing its FID in the few-step inference. We report the results of the eight-step inference in the COCO 2014, COCO 2017, and LAION dataset in Tab. 1, Tab. 2, and Tab. 3. Moreover, we use $p_{1}=5$, $p_{2}=1.2$ as the setting in the five- and six-step inference of generating 512 x 512 images and the six-step inference of generating 1024 x 1024 images. The result of the five- and six-step inference that synthesizes 512 x 512 images is also given in Tab. 1, Tab. 2, and Tab. 3. Moreover, the result of the 1024 x 1024 images is given in Tab. 4, Tab. 5, and Tab. 6. The free-U is used throughout the inference when we sample a 512 x 512 image, and it is used in the last three steps when we synthesize a 1024 x 1024 image. Moreover, the ODE solver, based on the analysis in Section 3, is DPM++2m for eight-step inference and DPM++1s for five-step inference. The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively. In the generation of 1024 x 1024 images in 6 steps, we use the SDE-Editing method mentioned in the paper[27]. Furthermore, all the synthesized images in Tab. 1 and Tab. 4 employs the random 10k COCO caption as a guidance prompt. In Tab. 2 and Tab. 5, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance. Furthermore, we use 10k captions in LAION for Tab. 3 and Tab. 6. We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2[43, 3, 19, 32]. Our approach outperforms state-of-the-art models for most of the datasets. However, some of the models[43] are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison. The $w$ in the Tab. 1 and Tab. 4 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
Solver FID CLIP-Score
(Our) custom DPM++1s(5) $w=7.5$ 19.18 30.92
(Our) custom DPM++1s(6) $w=7.5$ 18.54 31.16
(Our) custom DPM++2m(8) $w=7.5$ 15.92 31.32
(Our) custom DPM++1s(5) $w=5.5$ 17.54 31.0
(Our) custom DPM++1s(6) $w=5.5$ 17.04 31.16
(Our) custom DPM++2m(8) $w=5.5$ 15.7 31.07
Original DPM++2m(20) 17.3 30.97
Original DPM++2m(20) $w=5.5$ 18.68 30.97
Flash Diffusion(5) 17.34 30.45
Flash Diffusion(6) 17.94 30.59
Flash Diffusion(8) 19.03 30.37
AMED Pulgin(8) 19.07 31.15
SDXL turbo(4) 23.24 31.66
Table 1: Results of 512 x 512 image generation. COCO 2014
Solver FID CLIP-Score
(Our) custom DPM++1s(5) $w=5.5$ 24.22 31.06
(Our) custom DPM++1s(6) $w=5.5$ 23.24 31.12
(Our) custom DPM++2m(8) $w=5.5$ 22.40 31.23
(Our) custom DPM++1s(5) $w=7.5$ 25.58 30.93
(Our) custom DPM++1s(6) $w=7.5$ 24.71 31.13
(Our) custom DPM++2m(8) $w=7.5$ 22.35 31.35
Original DPM++2m(20) $w=7.5$ 25.97 31.21
Original DPM++2m(20) $w=5.5$ 23.75 31.10
Flash Diffusion(5) 23.56 30.64
Flash Diffusion(6) 24.16 30.60
Flash Diffusion(8) 25.35 30.55
AMED Pulgin(8) $w=7.5$ 25.50 31.17
SDXL turbo(4) 30.92 31.66
Table 2: Results of 512 x 512 image generation. COCO 2017
Solver FID CLIP Score
(Our) custom DPM++1s(5) $w=5.5$ 19.61 31.81
(Our) custom DPM++1s(6) $w=5.5$ 18.45 32.10
(Our) custom DPM++2m(8) $w=5.5$ 17.74 32.42
(Our) custom DPM++1s(5) $w=7.5$ 19.88 31.45
(Our) custom DPM++1s(6) $w=7.5$ 18.60 31.97
(Our) custom DPM++2m(8) $w=7.5$ 17.52 32.41
Original DPM++2m(20) $w=7.5$ 18.30 32.67
Original DPM++2m(20) $w=5.5$ 17.33 32.47
Flash Diffusion(5) 15.40 31.60
Flash Diffusion(6) 15.80 31.63
Flash Diffusion(8) 16.87 31.57
AMED Pulgin(8) $w=7.5$ 18.06 32.44
SDXL turbo(4) 22.25 33.02
Table 3: Results of 512 x 512 image generation. Laion
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 17.84 31.62
custom DPM++2m (8) $w=5.5$ 19.62 31.58
custom DPM++1s (6) $w=7.5$ 23 31.23
custom DPM++1s (6) $w=5.5$ 24.25 31.49
Flash DiffusionXL (6) 22.23 31.00
Flash DiffusionXL (8) 23.25 30.89
SDXL-Lightning (8) 21.07 31.11
DMD2 (4) 19.64 31.64
Table 4: Results of 1024 x 1024 image generation. COCO 2014
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 24.42 31.71
custom DPM++2m (8) $w=5.5$ 26.04 31.56
custom DPM++1s (6) $w=7.5$ 29.52 31.15
custom DPM++1s (6) $w=5.5$ 30.51 31.47
Flash DiffusionXL (6) 28.74 30.96
Flash DiffusionXL (8) 29.63 30.88
SDXL-Lightning (8) 27.89 31.08
DMD2 (4) 26.19 31.65
Table 5: Results of 1024 x 1024 image generation. COCO 2017
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 19.25 32.22
custom DPM++2m (8) $w=5.5$ 20.82 32.07
custom DPM++1s (6) $w=7.5$ 20.94 31.01
custom DPM++1s (6) $w=5.5$ 22.10 31.59
Flash DiffusionXL (6) 20.88 31.64
Flash DiffusionXL (8) 21.30 31.61
SDXL-Lightning (8) 20.45 32.50
DMD2 (4) 16.51 33.28
Table 6: Results of 1024 x 1024 image generation.Laion
We also calculate the PRD in a 6-step inference by sampling 10k captions from COCO 2014 and getting the corresponding image in the COCO dataset as a reference. PRD was first used to evaluate GAN. Considering two distributions $Q$ and $P$, the first one is the distribution of the generative images, and the second one is the distribution of the original images, the precision shows how much of $Q$ can be generated by $P$, and the recall exhibits how much $P$ can be generated by $Q$. For example, if $P$ is a subset of MNIST that only contains numbers ranging from 1 to 4, adding 5 and 6 to the generative distribution reduces the precision, while lacking modes like 3 and 4 will reduce the recall rate. Our experiment shows that the recall rate of the flash diffusion in the 6-step is less than that of us. The guidance scale of our sampler is 5.
Figure 5: The comparison of PRD between our 6-step inference sampler and the FLash diffusion.
5 Conclusion
In this study, we propose a plug-in method to enhance the sampling algorithm. By fully harnessing the potential capacity of the latent diffusion model—such as modifying the skip-connection at the right time, using the property of the decoder, which is a $\beta$-VAE trained with small KL-constraints, ignoring a small amount of noise in the final feature to reschedule the discretizing method, and choosing the proper ODE solver via truncation error analysis—we can sample a high-quality image with limited inference budget and avoid the need for additional training.
5.1 Limitation and the future work
A diffusion model is a robust generative model[4] that can prevent adversarial attacks. Most attacks that hack the latent diffusion model attack its VAE decoder. Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that mainly depends on the property of the VAE. Moreover, in future research, we explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
Acknowledgments
We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript.

### v2

@@
 Abstract
-The diffusion probability model is a state-of-the-art generative model that generates an image by applying a neural network iteratively.
-Moreover, this generation process is regarded as an algorithm solving a diffusion ordinary differential equation (ODE) or stochastic differential equation (SDE).
-Based on the analysis of the truncation error of the diffusion ODE and SDE, our study proposes a training-free algorithm that generates high-quality 512 x 512 and 1024 x 1024 images in eight steps, with flexible guidance scales.
-To the best of my knowledge, our algorithm is the first one that samples a 1024 x 1024 resolution image in 8 steps with an FID performance comparable to that of the latest distillation model, but without additional training.
-Meanwhile, our algorithm can also generate a 512 x 512 image in 8 steps, and its FID performance is better than the inference result using state-of-the-art ODE solver DPM++ 2m in 20 steps.
-The result of our algorithm in generating high-quality 512 x 512 images and 1024 x 1024 images with five-step and six-step inference is also comparable to the latest distillation model.
-Moreover, unlike most distillation algorithms, which achieve state-of-the-art FID performance by fixing the sampling guidance scale, and which sometimes cannot improve their performance by adding inference steps, our algorithm uses a flexible guidance scale on classifier-free guidance sampling.
-The increase in inference steps enhances its FID performance.
-Additionally, the algorithm can be considered a plug-in component compatible with most ODE solvers and latent diffusion models.
-Extensive experiments are performed using the COCO 2014, COCO 2017, and the LAION dataset.
-Specifically, we validate our eight-step image generation algorithm using the COCO 2014, COCO 2017, and LAION validation datasets with a 5.5 guidance scale and a 7.5 guidance scale, respectively.
-Furthermore, the FID performance of the image synthesis in 512 x 512 resolution with a 5.5 guidance scale is 15.7, 22.35, and 17.52, meaning it is comparable with the state-of-the-art ODE solver DPM++ in 20 steps, whose best FID performance is 17.3, 23.75, and 17.33, respectively.
-Further, it also outperforms the state-of-the-art AMED-plugin solver, whose FID performance is 19.07, 25.50, and 18.06.
-We also apply the algorithm in five-step inference without additional training, for which the best FID performance of our algorithm in COCO 2014, COCO 2017, and LAION is 19.18, 23.24, and 19.61, respectively, which is comparable to the performance of the state-of-the-art AMED Pulgin solver in eight steps, SDXL-turbo in four steps, and the state-of-the-art diffusion distillation model Flash Diffusion in five steps.
-Then, we validate our algorithm in synthesizing 1024 * 1024 images, whose FID performance in COCO 2014, COCO 2017, and LAION using eight-step inference is 17.84, 24.42, and 19.25, respectively.
-Thus, it outperforms the SDXL-lightning in eight steps, Flash DiffusionXL in eight steps, and DMD2 in four steps.
-Moreover, the FID performance of the six-step inference of our algorithm in the 1024 x 1024 image synthesis is 23, which only has a limited distance to the state-of-the-art distillation model mentioned above.
-We also use information theory to explain the advantage of our algorithm and why it achieves a strong FID performance.
+The diffusion model is a state-of-the-art generative model that samples images by applying a neural network iteratively.
+However, the original sampling algorithm requires substantial computation cost, and reducing the sampling step is a prevailing research area.
+To cope with this problem, one mainstream approach is to treat the sampling process as an algorithm that solves an ordinary differential equation (ODE).
+Our study proposes a training-free inference plugin compatible with most few-step ODE solvers.
+To the best of my knowledge, our algorithm is the first training-free algorithm to sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively.
+Based on analyses of the latent diffusion model’s structure, the diffusion ODE, and the Free-U mechanism, we explain why specific hyperparameter couplings improve stability and inference speed without retraining.
+Meanwhile, experimental results also reveal a new design space of the latent diffusion ODE solver.
+Additionally, we also analyze the difference between the original diffusion model and the diffusion distillation model via an information-theoretic study, which shows the reason why the few-step ODE solver designed for the diffusion model can outperform the training-based diffusion distillation algorithm in few-step inference.
+The tentative results of the experiment prove the mathematical analysis.
 1 Introduction
-The diffusion probability model (DM)[12], a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision.
-A typical diffusion model, such as stable diffusion (SD)[28], is inspired by the Brownian diffusion process in physics.
+The diffusion probability model (DM) proposed by Ho et al.
+(2020a), a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision.
+A typical diffusion model, such as stable diffusion (SD) (Rombach et al.
+(2022)), is inspired by the Brownian diffusion process in physics.
 It consists of two different processes: diffusion and reverse.
-The diffusion process adds small variants of Gaussian noise to the image step by step to facilitate its gradual conversion into pure Gaussian noise, while the reverse process uses the score matching[37] algorithm to train a neural network studying the gradient of the diffusion process by denoising noisy images[25].
-Then, the reverse of the diffusion process occurs.
-The diffusion process can be described by a diffusion stochastic differential equation (SDE) whose form is similar to the Brownian diffusion SDE.
-Moreover, the reverse process can be described by an ordinary differential equation (ODE) or an SDE[37].
-The ODE is equivalent to the SDE in the reverse process, which generates random variables step by step by transporting the standard Gaussian distribution to a target probability distribution by changing its geometric shape.
-However, it requires fewer inference steps to sample a high-quality image than solving the SDE, as it prevents involving a random effect during sampling.
-However, solving diffusion ODEs still requires a large amount of computational resources.
-Hence, solving this problem remains an active research area.
+The diffusion process adds small variants of Gaussian noise to the image step by step to convert an image into a pure Gaussian noise, while the reverse process uses the score matching algorithm (Song et al.
+(2020b)) to train a neural network studying the gradient of the diffusion process by denoising noisy images (Luo (2022)).
+Like the Brownian process in physics, the diffusion process can be described by a diffusion stochastic differential equation (SDE).
+Moreover, the research proposed by Song et al.
+(2020b) shows that there exists an ordinary differential equation (ODE) to describe the reverse process.
+And using the existing mathematical tool for ODE solving enhances the capability of the diffusion model to sample an image.
+However, compared to typical algorithms such as generative adversarial networks and variational autoencoders, solving diffusion ODEs still requires substantial computational resources.
+Hence, coping with this problem remains an active research area.
 Research related to it can be categorized into two classes.
-The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process[12, 24, 35, 45, 15, 21].
-The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and it reduces the inference steps by training a small distillation model[26, 36, 19, 44, 43, 31, 32, 16, 38, 22].
-The training-free method typically necessitates a 20-step inference[24], while algorithms requiring additional training can generate an image within four to eight steps.
-However, our results show that by selecting the proper hyperparameter of the ODE solver and adding a training free diffusion model decorator to exploit the capability of the latent diffusion model, the inference speed can exceed expectations.
-To that end, we propose a new inference method that can generate 512 x 512 images in eight steps without additional training, and its Frechet Inception Distance (FID)[10] performance surpasses that of the state-of-the-art ODE solver in 20 steps.
-Furthermore, when applying our algorithm in the five-step inference, it also outperforms the state-of-the-art distillation algorithm in the same number of steps.
-We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step inference exceeds most of the latest distillation models in eight steps.
-We perform extensive experiments using the COCO 2014[20], COCO 2017 [20], and LAION datasets[33], and exhibit the generation result in Fig 1.
+The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process (Ho et al.
+(2020a); Lu et al.
+(2025); Song et al.
+(2020a); Zhao et al.
+(2023); Karras et al.
+(2022); Liu et al.
+(2022)).
+The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and the algorithm reducing the inference steps is equivalent to train a small distillation model (Luo et al.
+; Song & Dhariwal (2023); Lin et al.
+(2024); Yin et al.
+(2024b; a); Salimans & Ho (2022); Sauer et al.
+(2024); Kim et al.
+(2023); Song et al.
+(2023); Liu et al.
+(2023)).
+The training-free method typically necessitates a 20-step inference, like the DPM++ solver proposed by Lu et al.
+(2025), while algorithms requiring additional training can generate an image within four to eight steps.
+However, our results show that by selecting the appropriate hyperparameter for the ODE solver and adding a training-free diffusion model decorator to leverage the latent diffusion model’s capabilities, the inference speed can exceed expectations.
+To that end, we propose a new inference method that generates 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al.
+(2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively.
+We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step and six-step inference exceeds most of the latest distillation models.
+We perform extensive experiments using the COCO 2014 (Lin et al.
+(2014)), COCO 2017 (Lin et al.
+(2014)), and LAION datasets (Schuhmann et al.
+(2022)), and exhibit the generation result in Fig 1.
 We also use information theory to explain why our algorithm achieves a strong FID result.
-Figure 1: The comparison between our algorithm and dpm++ solver in few step inference
+Figure 1: The comparison between our algorithm and UniPC solver in few step inference
 2 Related Work
 2.1 Diffusion ODE
-The image generation process that uses a diffusion model can be regarded as solving a special diffusion SDE or ODE.
+The image-generation process of the diffusion model can be deemed as solving a special diffusion ODE.
 A typical diffusion model has two processes: forward and reverse.
@@
 Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms.
-In this study, we use the variant preserve (VP) SDE[37] to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm[13].
+In this study, we use the variant preserve (VP) SDE proposed by Song et al.
+(2020b) to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm (Ho et al.
+(2020b)).
 Specifically, its choice is as follows:
-$f(t)=-\frac{1}{2}\beta(t),$ (2)
+$f(t)=-\frac{1}{2}\beta(t)=\frac{d\log{\alpha_{t}}}{dt},$ (2)
 $g(t)=\sqrt{\beta(t)},$ (3)
 $\beta(t)=\beta_{min}+t(\beta_{max}-\beta_{min}).$ (4)
-By solving the Eq.
-1 step by step with a given $x_{0}$, we obtain a noisy image $x_{t}$.
-Finally, when applying iteration in $T$ step, we obtain $x_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}I)$.
-Additionally, using the reparameterization algorithm, we can sample $x_{t}$ within one step[12].
-In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE[15]:
+Explicitly, in the diffusion process mentioned above, the variant of noise added at step $t$ is $\beta_{t}$.
+Explicilty, $\beta_{max}$ and $\beta_{min}$ can be determined arbitrarily if it can ensure $p(x_{T}|x_{0})$ is a standard Gaussian distribution.
+In the timestep $t$, the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the eq.˜5.
+$p(x_{t}|x_{0})=\mathcal{N}(\alpha(t)x_{0};\sigma^{2}(t)I).$ (5)
+Typically, to transform an image into a standard Gaussian noise, the DDPM algorithm requires $T=1000$ steps, and $\alpha_{T}$ is closed to the zero.
+Additionally, using the reparameterization algorithm, we can sample $x_{t}$ in a single step (Ho et al.
+(2020a)).
+In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE proposed by Karras et al. (2022):
 $\frac{dx}{dt}=\frac{\dot{s}(t)}{s(t)}x_{t}+s(t)^{2}\dot{\sigma}(t)\epsilon_{\theta}(x_{t},t,c),$ (6)
-where $c$ is the condition of input, like the caption that describes the image, $\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)s_{\theta}(x_{t},c)$, where $s_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)}$.
-Moreover, the score function normally used in VP SDE, it is trained by Eq.
-6
+where $c$ is the condition of input, like the caption that describes the image, and the $\epsilon_{\theta}$ is modeled by a neural network.
+The $\epsilon_{\theta}$ has a relation between the score function, which is described in the eq.˜7:
+$\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)score_{\theta}(x_{t},c),$
+$score_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)},$ (7)
+And the $p_{\theta}(x_{t},c)$ is the probability distribution of noisy images $x_{t}$ modeled implicitly by the neural network with parameters $\theta$.
+Moreover, the score function normally used in VP SDE is trained by eq.˜8
 […]
-and $q_{t}$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$.
-Specifically, the $s(t)$ and $\sigma(t)$ is defined by Eq. 7 and Eq. 8:
+and $q_{t}(x)$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$, and $q_{0}$ means the distribution of clean images.
+In the eq.˜6, the $s(t)$ and $\sigma(t)$ is defined by eq.˜9 and eq.˜10:
 $\sigma(t)=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1},$ (9)
@@
 where $\beta_{d}=\beta_{max}-\beta_{min}$.
-Using Eq.
-5 as a reverse ODE can help exploit the advantage of adapting the DDPM algorithm to the common diffusion framework proposed by Karras [15].
-However, the algorithm implementation should also consider whether a neural network can study this input–output pair if the input and output in different timesteps have an obvious gap in the variance when using this framework.
-So, to let the neural network study properly, we should formulate the equation as
-$\epsilon_{\theta}(x_{t},t,c)=c_{out}(\sigma_{t})F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$ (9)
-where $F_{\theta}$ is the real function that the neural network studies, $c_{in}$ scales the input to reduce its variance, $c_{out}$ scales the output to use the common framework proposed by Karras, and $c_{noise}(\sigma)$ transforms the variance $\sigma$ to the condition input $t$ of the neural network.
-For simplicity, the following section will use $D_{\theta}$ to denote the centrifuge of the possible denoised images predicted by the neural network when given a noisy image $x_{t}$:
-$D_{\theta}(x_{t};\sigma)=$ $c_{skip}(\sigma)x_{t}+$ (10)
-$c_{out}(\sigma)F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$
-where $c_{skip}(\sigma)=1$ in DDPM.
+Deeming the loss of the DDPM algorithm as a score function described by the eq.˜8 can help exploit the advantage of adapting the original diffusion model to the common diffusion ODE framework proposed by Karras et al.
+(2022).
 2.2 Diffusion ODE Solver
-Sampling an image from a diffusion model is equivalent to solving the Eq.
-5.
-Moreover, different solvers have been considered in previous research[15, 23, 24].
-The inference process approximates the trajectory of the reverse ODE by discretizing its path.
-The research on reverse ODE solvers attempts to precisely estimate the next step, thereby minimizing the truncation error caused by the discretization.
-As the culmination of the truncation error will affect the result of synthesis, minimizing the truncation error in each step is equivalent to subtracting the demand for an increased inference step that has caused precise discretization.
-Among those solvers, the DPM++ solver[24] is the latest state-of-the-art solver designed specifically for classifier-free guidance (CFG)[11] sampling.
-It predicts the next step via the following formula:
-$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}D_{\theta}(x_{t(\lambda)};\sigma_{t(\lambda)})d\lambda,$ (11)
-where $\lambda:=\log{a_{t}}-\log{\sigma_{t}}$ is a log signal-to-noise ratio (SNR) of the noisy image, while the function $(\lambda)$ transforms $\lambda$ to the corresponding $t$.
-We use the Taylor formula to obtain an approximation of the Eq. 11, and then there is a first-order DPM++1s solver:
-$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}e^{h}D_{\theta}(x_{s},\sigma_{s}),$ (12)
-in which $h=\lambda_{t}-\lambda_{s}$.
-Moreover, by using a multistep method, we obtain the second-order DPM++2m solver:
-$x_{t_{i}}=\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}x_{t_{i-1}}-\alpha_{t_{i}}(e^{-h_{i}}-1)D_{i},$ (13)
-where $D_{i}$ is as follows:
-$D_{i}$ $=(1+\frac{1}{2r_{i}})D_{\theta}(x_{t_{i-1}};\sigma_{t_{i-1}})$ (14)
-$-\frac{1}{2r_{i}}D_{\theta}(x_{t_{i-2}};\sigma_{t_{i-2}}),$
-and $r_{i}=\frac{h_{i-1}}{h_{i}}$.
+Sampling an image from a diffusion model is equivalent to solving the eq.˜6.
+Via the research from Lu et al.
+(2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the eq.˜11 :
+$x_{t}=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\epsilon_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (11)
+where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$.
+Sometimes, to cope with the high guidance scale on the classifier free guidance algorithm (CFG) (Ho & Salimans (2022)), researchers tend to replace eq.˜11 with the eq.˜12 to get a better result (Lu et al.
+(2025)):
+$x_{t}=\frac{\sigma({t})}{\sigma(s)}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}\hat{x}_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (12)
+where $\hat{x}_{\theta}$ is a clean feature predicted indirectly by the neural network.
+By using the Taylor Expand in the $\epsilon_{\theta}$ or $x_{\theta}$ (Lu et al. (2022; 2025)), we get the following eq.˜13:
+$x_{t}$ $=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\Sigma_{n=0}^{k-1}{\epsilon}^{(n)}_{\theta}(x_{t(\lambda)},\lambda)\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\frac{(\lambda-\lambda_{s})^{n}}{n!}d\lambda$
+$+\mathcal{O}(h_{i}^{k+1}),$ (13)
+However, calculating the neural networks’ deviation requires substantial computational resources.
+Researchers typically use the Runge-Kutta method (Lu et al.
+(2025); Zhao et al.
+(2023)) to approximate the deviation.
+Explicitly, the DPM++ 2s algorithm approximate $x_{t}$ via the following formula:
+$x_{t}$ $=\frac{\sigma({t})}{\sigma(s)}x_{s}-\alpha_{t}(e^{-h_{t}}-1)\hat{x}_{\theta}(x_{s},s)$
+$+\frac{1}{2r_{t}}*\alpha_{t}(e^{-h_{t}}-1)*(\hat{x}_{\theta}(\mu,m)-\hat{x}_{\theta}(x_{s},s)),$ (14)
+in which $r_{t}$ can be selected arbitarily between 0 and 1, and $h_{i}$, $m$ and $\mu_{i}$ is given below:
+$h_{t}$ $=\lambda_{t}-\lambda_{s},$ (15)
+$r_{t}$ $=\frac{\lambda_{m}-\lambda_{s}}{h_{i}},$ (16)
+$\mu_{t}$ $=\frac{\sigma_{m}}{\sigma_{s}}x_{s}-\alpha_{m}(e^{-r_{t}h_{t}}-1)\hat{x}_{\theta}(x_{s},s),$ (17)
+To get the first-order deviation of the $\epsilon_{\theta}$, the number of function evaluations (NFE) in DPM++ 2s is 2.
+While DPM++ 2m utilizes the latest cached output in the previous inference to calculate the approximation of the first-order deviation, and hence only requires 1 NFE, which is a typical multistep algorithm in ODE solving.
+Simultaneously, the UniPC and DPM V3 solver cached all of the previous output to correct and predict a high-order deviation.
+The eq.˜6 describes a continuous curved trajectory, and the ODE solver approximates it by discretizing the path with a series of lines.
+Research on reverse ODE solvers aims to precisely estimate the next step, thereby minimizing truncation error from pool discretization in the few-step inference.
+On the other hand, the discrete method proposed by Karras et al.
+(2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ to enhance the sample quality.
+Via the analysis of the Karras’ research, the $h_{t}$ should be large at the beginning of the inference, and then decrease monotonically.
 2.3 Diffusion Distillation
-The training-free method, which optimizes the truncation error in each step to minimize the demand for precise discretion.
-Unlike this method, the distillation algorithm[19, 32, 26, 36, 31] considers the whole inference process of applying a neural network iteratively as a large neural network.
-The distillation algorithm, which normally trains a small neural network, entails reducing its inference steps.
-However, training a neural network with high-quality content, a fast inference speed, and no additional parameters is difficult.
-Thus, the distillation algorithm, through the analysis of information theory[17] and experiments, can sometimes reduce the diversity of the generated images, thus reducing the FID performance.
+The training-free method optimizes the truncation error at each step and thus minimizes the need for additional inference steps for the precise discretization.
+Unlike this method, the distillation algorithm (Lin et al.
+(2024); Sauer et al.
+(2024); Luo et al.
+; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as an extensive neural network.
+The distillation algorithm, which typically trains a small neural network, reduces its inference steps.
+However, training a neural network with high-quality content, fast inference, and no additional parameters is difficult.
+Thus, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance.
 This approach may pose a potential risk in the future.
 AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules.
-Moreover, when people draw an image via their craft, they will sometimes be affected by the image they have previously watched.
-If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such generation will not be deemed plagiarism.
-However, generative models that use Internet images as training data are always denied by people who claim that if those models generate similar content in their dataset, it should be considered plagiarism.
-Furthermore, diversity is among the metrics that affect people’s impression of generative content.
+However, when a human draws an image via their craft, they will sometimes be affected by the image they have previously seen.
+If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such a generation will not be deemed plagiarism.
+Compared to this, generative models trained on Internet images are often criticized, with people claiming that if those models generate content similar to their dataset, it should be considered plagiarism.
+In such a situation, diversity is a metrics that affect people’s impression of generative content.
 2.4 Latent Diffusion Model
-Another method for reducing the calculation resources of the inference is to encode images into a low-dimensional latent space.
-Stable diffusion[28] performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image to use the attention mechanism in the U-Net[29] efficiently.
-One team’s experiment[28] shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism[40].
-In contrast, its perceptual information, which makes the image appear crisp, is strongly linked to local regions and can be modeled by a convolutional neural network (CNN)[18].
-Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder[2] to let the following inference ignore the perceptual detail.
-Then, it performs diffusion inference in latent space via a U-Net[29] with an attention mechanism to model the semantic information.
-Experiments show that compared with vector quantify VAE[41], the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.
+Another method for reducing the computational requirements of inference is to encode images into a low-dimensional latent space.
+Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image.
+One experiment from Rombach et al.
+(2022) shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism (Vaswani et al.
+(2017)).
+In contrast, its perceptual information, which provides the image with sufficient real details, is strongly tied to local regions and can be modeled by a convolutional neural network (CNN) (e.g.
+Li et al.
+(2021)).
+Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al.
+(2018)) to let the following inference ignore the perceptual detail.
+Then, it performs diffusion inference in latent space via a U-Net (Ronneberger et al.
+(2015)) with an attention mechanism to model the semantic information.
+Experiments show that compared with vector quantify VAE (Wu & Flierl (2020)), the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.
 3 Approach
 3.1 Algorithm Based on Truncation Error Analysis
-The basic logic of the diffusion ODE is as follows: When the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as detrimental as the non-noisy steps.
-These insights have been proved in a previous paper[15] via an experiment that discretizes the inference process by choosing a uniform variant gap in each discretized time step to obtain arrays: $\{\sigma_{i}=\frac{i}{N-1}\sigma_{max};i<64\}$ and $\{t_{i};t(\sigma_{i})\}$.
-Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$.
-The experimental results show that the truncation error increases monotonically as the start variant decreases.
-[…]
-Based on this observation, one researcher [15] proposes a following discrete method:
-$\sigma_{i}=(\sigma_{min}^{\frac{1}{p}}+\frac{i}{N-1}(\sigma_{min}^{\frac{1}{p}}-\sigma_{max}^{\frac{1}{p}}))^{p}.$ (15)
-Eq.
-15 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part.
-Moreover, according to the observation from the paper[17], if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget.
-Considering the U-Net has 860 million parameters, and the variant of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical.
-Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
-Based on that observation, one study [17] proposes another discrete method:
-$t_{i}=(t_{min}^{\frac{1}{p}}+\frac{i}{N-1}(t_{min}^{\frac{1}{p}}-t_{max}^{\frac{1}{p}}))^{p}.$ (16)
-When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part is larger than that in Karras’s original method.
-Furthermore, we should consider the meaning of the score function by rewriting the Eq. 6 as follows:
+To design a training-free, few-step inference algorithm, we should first consider the meaning of the original diffusion models’ output, the score function.
+And then compare it with the distillation model’s output.
+By rewriting the eq.˜8 as follows:
 $\mathcal{L}_{\theta}=\frac{E_{x_{0}}[q(x_{t}|x_{0})||s_{\theta}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}]}{E_{x_{0}}[q(x_{t}|x_{0})]},$ (18)
-which means the score function is pointing to the average of all possible images that can cause this noisy image.
-Naturally, one can surmise that $\nabla_{x}\log q(x_{t})\approx\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage.
+We find out that the score function is pointing to the average of all possible images that can cause this noisy image.
+In an extreme few-step inference case, like one or two NFE inference, the final output will be blurred, since in that step, the score function only points to the centroid of images.
+Meanwhile, the distillation algorithm, like rectify flow proposed by Liu (2022), will tend to modify the output and let it point to a special sample.
+Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, one can presume that the score function $\nabla_{x}\log q(x_{t})$ can be approximated by the conditional score function $\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage, since the possible images are very limited.
 In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
-Based on the aforementioned idea and the experiment results in Section 4, we propose the following discrete method that aims to correct the output in the less noisy stage:
-$t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}},$ (18)
-where $\sigma_{stop}$ is defined via the following formula:
-$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{N+stop+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (19)
-which is not zero, as the decode $\beta$-VAE trained with a small KL-term constraint can handle a small amount of noise, and this setting can reduce the one-step inference budget.
-The visual different between our discrete method and the method proposed by the Karras’ paper is in Fig.
-2
-Figure 2: The comparison between our time schedule and the time schedule proposed by Karras
-Additionally, consider the Eq.
-17; the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, which can be achieved by modifying the skip connection and the backbone feature of the U-Net.
-Previous study[34] adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability.
-Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the special part of the noisy backbone feature map that lower than a particular frequency.
-The decorator only applies to the first two skip connections, and the scaler affected the backbone feature is $b_{1}$ and $b_{2}$, while the scaler affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
-Also, consider that the normal ResNet[9] can be viewed as an ODE[5] by using the following formula:
+Since the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, one of the natural method is to enhance the denoise capability of U-Net, which can be achieved by modifying the skip connection and the backbone feature of the U-Net.
+Previous study (Si et al.
+(2024)) adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability.
+Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the low-frequency part of the noisy backbone feature map.
+The decorator only applies to the first two skip connections.
+The scaler that affects the backbone feature is $b_{1}$ and $b_{2}$, while the scaler that affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
+However, we don’t know which step could be changed simply from the score function to the conditional score function, and one of the vanilla ways is to apply free-U at all time steps.
+This method has been seen as helpful to slightly modify the score function and make it more similar to the distillation’s output.
+The previous research from Ma et al.
+(2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers, since the skip connection itself will abate the result in the block below that layer, while the few-step inference tends to excute a huge transportation from a standard Gaussian distribution to a complex probability distribution, and hence require the skip connection reduce its’ effect that abate the block below it.
+However, via the mathematical analysis of the diffusion ODE and the experimental result.
+This vanilla method in the extreme few-step inference of high-resolution images, such as 6 NFE, would not work properly.
+Explicitly, consider that the normal ResNet (He et al. (2016)) can be viewed as an ODE (Chen et al. (2018)) by using the following formula:
 $h_{t+1}=h_{t}+f_{\theta}(h_{t},t),$ (19)
@@
 The U-Net, as a black box ODE solver, can be regarded as another ODE.
-Thus, modifying the skip connection and scale backbone feature will further increase its output and the moving distance of the diffusion reverse ODE, which let the start step of using U-Net decorator should be parameterized in 1024 x 1024 image synthesis, as it contributes to the increase of moving distance in the noisy part.
-This behavior causes the global truncation error in solving ODE, which may let the variables be out of the distribution (OOD).
-We also visualize the decoded feature in the trajectory of different samplers in Fig.
-3.
-Figure 3: The comparison between our sampler and the original DPM++1s sampler.
+Thus, modifying the skip connection and scale backbone feature will further increase the norm of the output, increasing the moving distance of the diffusion reverse ODE.
+In the situation of the few-step inference, it will be a serious problems, since the weakness of the poor discretization method will be amplified by this behaviour.
+The related mathematic details will be mentioned below.
+We also visualize the decoded features along the trajectories of different samplers in fig.˜2.
+Figure 2: The comparison between our custom DPM++1s sampler and the original DPM++1s sampler.
 The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler.
 The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
-We then visualize the decode trajectory of using different discrete methods by applying the same sampler DPM++ 1s sampler in Fig.
-5.
-The normal time schedule using Free-U will generate a trajectory similar to the method using Karras’ time schedule without using Free-U, which exhibits the evidence of our previous claim.
-Figure 4: The comparison between different discrete method when using the same DPM++ 1s Sampler in a few-step inference.
-The first line exhibit the trajectory of the normal Karras discrete method.
-The second one exhibit the trajectory of the Karras discrete method + Free-U.
-The third exhibite the normal discrete method with normal U-Net, and the final one exhibit the normal discrete method with Free-U.
-And the final line is similiar to the first line.
-Our algorithm can be described by the Algorithm 1
-Data: number of inference step $N$, initiated noise $\epsilon$, stop step in karras schedule $t_{stop}$, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$
+To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by applying the same sampler DPM++ 1s sampler in fig.˜3.
+The standard time schedule using Free-U generates a trajectory similar to that of Karras et al.
+(2022) time schedule without Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference, which determines the image layout composition.
+Figure 3: The comparison between different discrete methods when using the same DPM++ 1s Sampler in a few-step inference.
+The first line exhibits the trajectory of the normal Karras discrete method.
+The second one exhibits the trajectory of the Karras discrete method + Free-U.
+The third exhibits the normal discrete method with a normal U-Net, and the final one exhibits the normal discrete method with Free-U.
+And the final line is similar to the first line.
+Another method is to parameterize the start step of using the U-Net decorator in an existing solver.
+However, experimental results show that this Vallina method doesn’t work either.
+We visualize the 6-step inference result combining the free-U and UniPC solver in fig.˜4.
+Figure 4: The result that starts using free-U in the different steps of the UniPC solver.
+To address black-box ODE-solving problems, we should identify the characteristics of the diffusion ODE.
+The basic logic of the ODE solver is to discretize a continuous curve trajectory into a series of lines.
+And to achieve the best performance, the ODE solver should determine the length of each discrete line for different ODEs, such as the Fehlberg method (see e.g.
+Atkinson et al.
+(2009)) that adjusts the discrete length based on the error per unit stepsize.
+For the diffusion ODE, when the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as fatal as the non-noisy steps.
+These insights have been proved by Karras et al.
+(2022) via an experiment that discretizes the inference process with a uniform gap between each timestep.
+And it obtains an array of steps $\{t_{i}\}$.
+Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$.
+The experimental results show that the truncation error increases monotonically as the noise variance add in the input feature at the time step $t_{i}$ decreases.
+[…]
+Based on this observation, Karras et al. (2022) proposes a following discrete method:
+$\sigma_{i}=(\sigma_{min}^{\frac{1}{p}}+\frac{i}{N-1}(\sigma_{min}^{\frac{1}{p}}-\sigma_{max}^{\frac{1}{p}}))^{p}.$ (21)
+eq.˜21 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part.
+Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget.
+Considering the U-Net has 860 million parameters, and the variance of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical.
+Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
+Based on that observation, Li & Zhang (2025) proposes another discrete method:
+$t_{i}=(t_{min}^{\frac{1}{p}}+\frac{i}{N-1}(t_{min}^{\frac{1}{p}}-t_{max}^{\frac{1}{p}}))^{p}.$ (22)
+When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method.
+In such a situation, $t_{i}-t_{i-1}$ will increase in the noisier part.
+The eq.˜4 shows that $t_{i}$ is proportional to the variance of noise added in $x_{t}$, which can be deemed as the speed of movement at timestep $t_{i}$.
+The increase of $t_{i}-t_{i-1}$ in the noisy part aligns with the rise in speed, while the rise of $\sigma_{i}-\sigma_{i-1}$ aligns with the increment of the moving distance, hence it can cope with the problem of the original Karras schedule mentioned above.
+However, both of those discrete methods have been commonly used in all kinds of images.
+Our research shows that, even with the same time schedule and UniPC solver, perceptual quality in the few-step inference remains substantially different across image resolution and image type.
+Hence, the design space of hyperparameters for different ODE solvers has not yet been totally exploited.
+We find a discrete method, which changes the eq.˜23, a formula with a special design reason mentioned later, by adding one more step to get clean denoised images.
+And use it to generate images at different resolutions and of different types in a few-step inference.
+We find that the result in fig.˜5 has a noticeable gap in perceptual quality.
+Figure 5: The comparison between the different resolution images and different type images.
+The first line contains 512 x 512 images generated by Stable Diffusion 1.5, the second line contains 1024 x 1024 images generated by Dreamshaper XL.
+The last lines contain high-resolution ACGN images generated by Illustrij Evo.
+Based on the experimental results, designing a customized training-free algorithm for different image types requires exploring the design space of the ODE solver for the latent diffusion model.
+Firstly, the decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget.
+Although the latent diffusion model theoretically does not require using the KL term to train an autoencoder, this optimization trick is very prevalent in the training of the diffusion model.
+The latest representation autoencoders (RAE) also add noise during the training to augment the decoder Zheng et al.
+(2025).
+The experimental results show that the decoder’s capacity to handle noisy features varies with image resolution.
+In the fig.˜6, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem.
+And the third line output feature with less noise, and has a better results.
+Figure 6: The comparison between the different resolution images generated by the same time schedule and ODE solver utilizing the same $\beta$-VAE.
+Based on the aforementioned idea and the experiment results in Section 4, we first involve the following discrete method that aims to correct the output in the less noisy stage:
+$t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}},$ (23)
+where $\sigma_{stop}$ is defined via the following formula:
+$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{M}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (24)
+which could be nonzero to save the one-step inference budget.
+By applying this special time schedule, the problems of using Free-U while preventing an increase in moving distance could be addressed, since we could only apply Free-U at timesteps that move a short distance.
+By leveraging the robustness of the $\beta$-VAE, the total moving distance has also been constrained.
+The visual difference between our discrete method and the method proposed by the Karras’ paper is in fig.˜7
+Figure 7: The comparison between our time schedule and the time schedule proposed by Karras
+Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) during inference.
+Our algorithm can be described by the algorithm˜1
+Data: number of inference step $N$, initiated noise $\epsilon$, stop step in Karras schedule $t_{stop}$, the total step $M$ of the original discrete method, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$, a bool value afsafs to determine whether using afs, and an analytical first step solver without using neural network $F$
 Result: final result image
 1 $\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$;
-2 $ts=\{t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}};0\leq i<N\}$ $x=\epsilon$ while $t$ in tsts do
-3 if $t$ equal to $t_{aug}$ then
-4    $freeU(F_{\theta})$;
-5
+2
+[…]
+4if afsafs then
+5 $ts.insert((\frac{t_{0}+t_{1}}{2}),at:1)$
 6 end if
-7 $x=F_{\theta}(x)$
+7
+[…]
+9while $t$ in tsts do
+10 if $t$ equal to $t_{aug}$ then $freeU(F_{\theta})$;
+11   ;
+12 if afsafs and $t$ equal to 999999 then $x=F(x)$ ;
+13 else $x=F_{\theta}(x)$ ;
+14
 15 end while
@@
 3.2 Information Theory Analysis of the Different Algorithms
-A common point of confusion in deep learning at an early age is how neural networks with so many parameters still have generalization ability[1].
-The normal statistical model uses an extremely limited number of parameters to ensure that it can concisely describe the phenomenon and capture the essence of the observed data.
-For this problem, the study theory[8] claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions.
+The experimental results in the section˜4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance.
+Given that all images are generated from the same prompts, and the distillation algorithm changes the score function from pointing to the centroid of all possible images to a special image, we assume the distillation model limits the diversity of images’ low-frequency components, which is caused by modification of the score function at the start stage of the inference.
+Explicitly, some images with totally different semantics could have a similar, fuzzy structure with the same color block in a similar position.
+And the limitation on the possible fuzzy structure of images may also cause diversity problems.
+And we use information theory to explain why applying Free-U only in the final stage of the inference helps address this problem.
+To introduce the relationship between information theory and machine learning, we should first review the historical confusion about machine learning’s generalization capability.
+A common point of confusion in deep learning at an early stage is how neural networks with so many parameters can still generalize (Breiman (2018)).
+The standard statistical model uses a limited number of parameters to concisely describe the phenomenon and capture the essence of the observed data, whereas deep learning is fundamentally different.
+For this problem, the study theory (Harman & Kulkarni (2012)) claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions.
 Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{N}$.
-According to the information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can gain access to is $2^{H(D)}$[42].
+According to information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can access is $2^{H(D)}$ (Yeung (2012)).
+For example, in a high-dimensional standard Gaussian distribution, a pure black image, although with the highest possible density, is not in the typical set.
 Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
 Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$.
-The information bottleneck theory[39] shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ project to the $Z$.
-Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses.
-As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) [7], which face mode collapse[6] that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint [14].
-One study[17] shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information.
-The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE)[2], in which, when $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bounce of the mutual information constraint will be abated, and there will be an increase in the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$.
-From this perspective, the diffusion model can be deemed as a hierarchy VAE[25], while progressive distillation[31] increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$.
+The information bottleneck theory proposed by Tishby et al.
+(2000) shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ projected to the $Z$.
+Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses by limiting the possible middle feature (Tishby et al.
+(2000)).
+As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) (Goodfellow et al.
+(2020)), which face mode collapse (Creswell et al.
+(2018)) that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint Jeon et al.
+(2021).
+One study (Li & Zhang (2025)) shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information.
+The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE).
+When $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bound of the mutual information constraint will be abated.
+Meanwhile, the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$ will increase.
+From this perspective, the diffusion model can be deemed as a hierarchy VAE (Luo (2022)), while progressive distillation proposed by Salimans & Ho (2022) increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$.
 Where $X_{\theta}$ is a noisy feature output by solving ODE.
-Hence, the PSNR also increases.
-A recent example of diffusion distillation is SDXL-lightning[19], which adds GAN loss to the progressive distillation training, and its FID remains similar to normal progressive distillation.
+Hence, the PSNR of the middle output feature also increases.
 However, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will also affect the mutual information, even if the output does not have an increased PSNR.
 We provide further analysis regarding this.
-The diffusion ODE can be regarded as a Markov process $X_{\theta}(T)-X_{\theta}(T-1)-...-X_{\theta}(t)-...X_{\theta}(0)$, in which $X_{\theta}(t)$ follows the input of the diffusion model predicted by the U-Net, and we have the following Eq.
-23:
-$I(X_{\theta}(0),X_{\theta}(t))=H(X_{\theta}(0))-H(X_{\theta}(0)|X_{\theta}(t))$ (22)
-$\leq H(X(0))-E_{X_{\theta}(t)}[p(X_{\theta}(0)|X_{\theta}(t),$
-$*\log p_{\theta}(X_{\theta}(0)|X_{\theta}(t)))],$ (23)
-in which $H(X(0))$ is the real $X(0)$ sample from the dataset.
-In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$ while enhancing the log likelihood of the final synthesis.
-Moreover, Eq.
-23 shows that if we modify the $X_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase.
+The diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following eq.˜25:
+$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))$
+$\leq H(X(0))-E_{Z_{\theta}(t)}[p(Z_{\theta}(0)|Z_{\theta}(t))$
+$*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))],$ (25)
+in which $H(X(0))$ is the entropy of the real $X(0)$ sample from the dataset.
+In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$, enhancing the log likelihood of the $Y$.
+And in the distillation algorithm, eq.˜25 shows that if we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase.
 Most distillation methods that do not increase the PSNR of the output will still face this problem.
-We evaluate the Precision and Recall of Distribution (PRD)[30] of flash diffusion and our sampler to further validate this claim.
+We evaluate the Precision and Recall of Distribution (PRD) (Sajjadi et al.
+(2018)) of flash diffusion and our sampler to further validate this claim.
 4 Experiments
+4.1 Ablation Study
+We conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampling from the COCO 2017 dataset.
+Explicitly, for the 8-step inference, we employ $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ as the discrete setting.
+And for the 6-step inference, the discrete setting is $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$.
+Then, we try the following setting: 1.
+apply Free-U on every step of inference, 2.
+apply Free-U only on the last few steps of inference, 3.
+Don’t apply Free-U.
+The experimental result is in the table˜1.
+The $w$ in the table˜1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
+Meanwhile, for the 8-step UniPC solver, we use a 2nd-order multistep algorithm, whereas the 6-step solver is 1st-order solver.
+Solver FID CLIP-Score
+custom UniPC(6) apply Free-U on the every steps $w=5.5$ 28.88 30.96
+custom UniPC(8) apply Free-U on the every steps $w=5.5$ 33.70 30.50
+custom UniPC(6) don’t apply Free-U $w=5.5$ 27.13 31.49
+custom UniPC(8) don’t apply Free-U $w=5.5$ 25.47 31.39
+custom UniPC(6) start apply Free-U at 4th step $w=5.5$ 25.20 31.42
+custom UniPC(8) start apply Free-U at 6th step $w=5.5$ 24.38 31.38
+custom UniPC(6) apply Free-U on the every steps $w=7.5$ 39.62 30.33
+custom UniPC(8) apply Free-U on the every steps $w=7.5$ 54.26 29.63
+custom UniPC(6) don’t apply Free-U $w=7.5$ 25.34 31.53
+custom UniPC(8) don’t apply Free-U $w=7.5$ 24.61 31.22
+custom UniPC(6) start apply Free-U at 4th step $w=7.5$ 25.29 31.44
+custom UniPC(8) start apply Free-U at 6th step $w=7.5$ 24.58 31.21
+Table 1: Results of applying Free-U in different steps of inference.
 4.2 Performance
-In this study, we employ $p_{1}=7$, $p_{2}=1.2$, $stop=3$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images.
-For the eight-step inference, as the precise discretization method, we can completely avoid using free-U, which can enhance the aesthetic of images while reducing its FID in the few-step inference.
-We report the results of the eight-step inference in the COCO 2014, COCO 2017, and LAION dataset in Tab.
-1, Tab.
-2, and Tab.
-3.
-Moreover, we use $p_{1}=5$, $p_{2}=1.2$ as the setting in the five- and six-step inference of generating 512 x 512 images and the six-step inference of generating 1024 x 1024 images.
-The result of the five- and six-step inference that synthesizes 512 x 512 images is also given in Tab.
-1, Tab.
-2, and Tab.
-3.
-Moreover, the result of the 1024 x 1024 images is given in Tab.
-4, Tab.
-5, and Tab.
-6.
-The free-U is used throughout the inference when we sample a 512 x 512 image, and it is used in the last three steps when we synthesize a 1024 x 1024 image.
-Moreover, the ODE solver, based on the analysis in Section 3, is DPM++2m for eight-step inference and DPM++1s for five-step inference.
+In this study, we employ our algorithm on the DPM++ 2m solver and UniPC solver.
+We employ $p_{1}=7$, $p_{2}=1.2$, $stop=8$, $M=12$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images via DPM++ 2m.
+Additionally, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images.
+And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$.
+Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$.
+And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$.
+Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
+For the eight-step inference that uses DPM++ 2m to generate 512x512-resolution images, we avoid using free-U.
+And for the customized DPM++ 2m solver that generates 1024x1024-resolution images, the $t_{aug}$ apply free-U is 8.
+For all UniPC solver customized to generate 512x512 resolution images, $t_{aug}=4$.
+And $t_{aug}=6$, $t_{aug}=4$ for 6-step UniPC solver and 8-step UniPC solver customeized for 1024 x 1024-resolution images, respectively.
+And we only use AFS on the UniPC solver customized for the 1024 x 1024-resolution image.
+We don’t guarantee we provide the best hyperparameters; we just ensure we provide hyperparameters that can sample images with FID performance comparable to that of the inference result from the 20-step DPM++ solver and the SOTA distillation algorithm.
+We report the results of the inference that generate 512 x 512-resolution images in the COCO 2014, COCO 2017, and LAION dataset in table˜2, table˜3, and table˜4.
+Moreover, the result of the 1024 x 1024 images is given in table˜5, table˜6, and table˜7.
+Explicitly, the UniPC solver is a 2nd-order multistep solver for eight-step inference and a 1st-order multistep solver for five and six-step inference.
 The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively.
-In the generation of 1024 x 1024 images in 6 steps, we use the SDE-Editing method mentioned in the paper[27].
-Furthermore, all the synthesized images in Tab.
-1 and Tab.
-4 employs the random 10k COCO caption as a guidance prompt.
-In Tab.
-2 and Tab.
-5, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance.
-Furthermore, we use 10k captions in LAION for Tab.
-3 and Tab.
-6.
-We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2[43, 3, 19, 32].
+Furthermore, all the synthesized images in table˜2 and table˜5 employs the random 10k COCO caption as a guidance prompt.
+In table˜3 and table˜6, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance.
+Furthermore, we use 10k captions in LAION for table˜4 and table˜7.
+We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2 (Yin et al.
+(2024a); Chadebec et al.
+(2025); Lin et al.
+(2024); Sauer et al.
+(2024)).
 Our approach outperforms state-of-the-art models for most of the datasets.
-However, some of the models[43] are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison.
-The $w$ in the Tab.
-1 and Tab.
-4 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
+However, some of the models (Yin et al.
+(2024a); Chadebec et al.
+(2025); Lin et al.
+(2024)) are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison.
+The $w$ in the table˜2 and table˜5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
 The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
@@
 (Our) custom DPM++2m(8) $w=5.5$ 15.7 31.07
-Original DPM++2m(20) 17.3 30.97
+(Our) custom UniPC(5) $w=7.5$ 19.35 30.63
+(Our) custom UniPC(6) $w=7.5$ 16.41 30.97
+(Our) custom UniPC(8) $w=7.5$ 16.00 31.15
+(Our) custom UniPC(5) $w=5.5$ 15.19 30.93
+(Our) custom UniPC(6) $w=5.5$ 14.72 31.08
+(Our) custom UniPC(8) $w=5.5$ 15.26 31.12
+Original DPM++2m(20) $w=7.5$ 17.3 30.97
 Original DPM++2m(20) $w=5.5$ 18.68 30.97
@@
 (Our) custom DPM++2m(8) $w=7.5$ 22.35 31.35
+(Our) custom UniPC(5) $w=7.5$ 22.70 30.99
+(Our) custom UniPC(6) $w=7.5$ 24.02 31.04
+(Our) custom UniPC(8) $w=7.5$ 24.55 30.99
+(Our) custom UniPC(5) $w=5.5$ 21.31 30.92
+(Our) custom UniPC(6) $w=5.5$ 20.82 31.04
+(Our) custom UniPC(8) $w=5.5$ 21.71 31.07
 Original DPM++2m(20) $w=7.5$ 25.97 31.21
@@
 (Our) custom DPM++2m(8) $w=7.5$ 17.52 32.41
+(Our) custom UniPC(5) $w=7.5$ 20.43 30.87
+(Our) custom UniPC(6) $w=7.5$ 17.56 31.63
+(Our) custom UniPC(8) $w=7.5$ 17.38 32.08
+(Our) custom UniPC(5) $w=5.5$ 17.44 31.73
+(Our) custom UniPC(6) $w=5.5$ 16.84 32.08
+(Our) custom UniPC(8) $w=5.5$ 17.32 32.26
 Original DPM++2m(20) $w=7.5$ 18.30 32.67
@@
 custom DPM++2m (8) $w=5.5$ 19.62 31.58
-custom DPM++1s (6) $w=7.5$ 23 31.23
-custom DPM++1s (6) $w=5.5$ 24.25 31.49
+custom UniPC (8) $w=7.5$ 18.18 31.23
+custom UniPC (8) $w=5.5$ 17.87 31.40
+custom UniPC (6) $w=7.5$ 18.80 31.42
+custom UniPC (6) $w=5.5$ 19.07 31.45
 Flash DiffusionXL (6) 22.23 31.00
@@
 custom DPM++2m (8) $w=5.5$ 26.04 31.56
-custom DPM++1s (6) $w=7.5$ 29.52 31.15
-custom DPM++1s (6) $w=5.5$ 30.51 31.47
+custom UniPC (8) $w=7.5$ 24.58 31.21
+custom UniPC (8) $w=5.5$ 24.38 31.38
+custom UniPC (6) $w=7.5$ 25.29 31.44
+custom UniPC (6) $w=5.5$ 25.20 31.42
 Flash DiffusionXL (6) 28.74 30.96
@@
 custom DPM++2m (8) $w=5.5$ 20.82 32.07
-custom DPM++1s (6) $w=7.5$ 20.94 31.01
-custom DPM++1s (6) $w=5.5$ 22.10 31.59
+custom UniPC (8) $w=7.5$ 21.59 31.73
+custom UniPC (8) $w=5.5$ 20.41 32.06
+custom UniPC (6) $w=7.5$ 22.51 31.82
+custom UniPC (6) $w=5.5$ 22.33 31.87
 Flash DiffusionXL (6) 20.88 31.64
@@
 5.1 Limitation and the future work
-A diffusion model is a robust generative model[4] that can prevent adversarial attacks.
+Firstly, a diffusion model is a robust generative model (Chen et al.
+(2024)) that can prevent adversarial attacks.
 Most attacks that hack the latent diffusion model attack its VAE decoder.
-Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that mainly depends on the property of the VAE.
-Moreover, in future research, we explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
-Acknowledgments
-We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript.
+Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that largely depends on the property of the VAE.
+Moreover, in future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
+Last, the hyperparameters of the ACGN model need further exploration.

### v3

@@
-Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model
+F-scheduler: illuminating the free-lunch design space for fast sampling of diffusion models
 Abstract
-The diffusion model is a state-of-the-art generative model that samples images by applying a neural network iteratively.
-However, the original sampling algorithm requires substantial computation cost, and reducing the sampling step is a prevailing research area.
-To cope with this problem, one mainstream approach is to treat the sampling process as an algorithm that solves an ordinary differential equation (ODE).
-Our study proposes a training-free inference plugin compatible with most few-step ODE solvers.
-To the best of my knowledge, our algorithm is the first training-free algorithm to sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively.
-Based on analyses of the latent diffusion model’s structure, the diffusion ODE, and the Free-U mechanism, we explain why specific hyperparameter couplings improve stability and inference speed without retraining.
-Meanwhile, experimental results also reveal a new design space of the latent diffusion ODE solver.
-Additionally, we also analyze the difference between the original diffusion model and the diffusion distillation model via an information-theoretic study, which shows the reason why the few-step ODE solver designed for the diffusion model can outperform the training-based diffusion distillation algorithm in few-step inference.
-The tentative results of the experiment prove the mathematical analysis.
+Diffusion models are the state-of-the-art generative models for high-resolution images, but sampling from pretrained models is computationally expensive, motivating interest in fast sampling.
+Although Free-U Net is a training-free enhancement for improving image quality, we find it ineffective under few-step (<10<10) sampling.
+We analyze the discrete diffusion ODE and propose F-scheduler, a scheduler designed for ODE solvers with Free-U Net.
+Our proposed scheduler consists of a special time schedule that does not fully denoise the feature to enable the use of the KL-term in the $\beta$-VAE decoder, and the schedule of a proper inference stage for modifying the U-Net skip-connection via Free-U Net.
+Via information theory, we provide insights into how the better scheduled ODE solvers for the diffusion model can outperform the training-based diffusion distillation model.
+The newly proposed scheduler is compatible with most of the few-step ODE solvers and can sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps when it applies to DPM++ 2m and UniPC, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively.
+Machine Learning, ICML
 1 Introduction
-The diffusion probability model (DM) proposed by Ho et al.
-(2020a), a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision.
-A typical diffusion model, such as stable diffusion (SD) (Rombach et al.
-(2022)), is inspired by the Brownian diffusion process in physics.
-It consists of two different processes: diffusion and reverse.
-The diffusion process adds small variants of Gaussian noise to the image step by step to convert an image into a pure Gaussian noise, while the reverse process uses the score matching algorithm (Song et al.
-(2020b)) to train a neural network studying the gradient of the diffusion process by denoising noisy images (Luo (2022)).
-Like the Brownian process in physics, the diffusion process can be described by a diffusion stochastic differential equation (SDE).
-Moreover, the research proposed by Song et al.
-(2020b) shows that there exists an ordinary differential equation (ODE) to describe the reverse process.
-And using the existing mathematical tool for ODE solving enhances the capability of the diffusion model to sample an image.
-However, compared to typical algorithms such as generative adversarial networks and variational autoencoders, solving diffusion ODEs still requires substantial computational resources.
-Hence, coping with this problem remains an active research area.
+The diffusion probability model (DPM) proposed by Ho et al.
+(2020) is a state-of-the-art image generation model.
+However, compared to typical algorithms (Goodfellow et al., 2020; Kingma & Welling, 2013), sampling from DPMs requires multiple neural network evaluations and incurs substantial computational cost.
+Hence, accelerating the sampling from DPMs remains an active research area.
 Research related to it can be categorized into two classes.
 The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process (Ho et al.
-(2020a); Lu et al.
+(2020); Lu et al.
 (2025); Song et al.
-(2020a); Zhao et al.
+(2020); Zhao et al.
 (2023); Karras et al.
@@
 (2022)).
-The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and the algorithm reducing the inference steps is equivalent to train a small distillation model (Luo et al.
+The second class deems the whole inference process as a large neural network, and reduces the inference steps by training a small distillation model (Luo et al.
 ; Song & Dhariwal (2023); Lin et al.
 (2024); Yin et al.
-(2024b; a); Salimans & Ho (2022); Sauer et al.
+(2024b, a); Salimans & Ho (2022); Sauer et al.
 (2024); Kim et al.
@@
 (2023); Liu et al.
-(2023)).
-The training-free method typically necessitates a 20-step inference, like the DPM++ solver proposed by Lu et al.
-(2025), while algorithms requiring additional training can generate an image within four to eight steps.
-However, our results show that by selecting the appropriate hyperparameter for the ODE solver and adding a training-free diffusion model decorator to leverage the latent diffusion model’s capabilities, the inference speed can exceed expectations.
-To that end, we propose a new inference method that generates 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al.
-(2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively.
+(2023)), which can generate images in four to eight steps.
+Most training-free ODE solvers are not tailored to diffusion models, except the method of Karras et al.
+(2022), which shows that a uniform discretizing schedule causes increasing truncation error in few-step inference and proposes a non-uniform schedule with larger steps in high-noise regions and smaller steps in low-noise regions.
+However, this method is not customized for the latent diffusion model.
+The latent diffusion model itself includes $\beta$-VAE and skip connections, both of which can be additionally boosted by a better scheduler in the design space.
+The $\beta$-VAE trained with a little noise can also reduce the inference budget by preventing the total denoising of images.
+However, adding same noise to the latent feature at different resolutions also has distinct effects(Hoogeboom et al., 2023), which necessitates the investigation of the design space.
+Furthermore, we find that the U-Net decorator, which aims to change the skip connection to enhance the denoising process, has the side effect of mimicking the distillation model(Ma et al., 2024) and changing the discretization result.
+We use ODE theory to fix the unwilling discretization effect, and use information theory to explain why the distillation model has the diversity problem, which provides an explanation of the phenomenon proposed by Bertrand et al.
+(2025), and proves the advantage of exploiting the design space of using a skip-connection decorator.
+To that end, by exploring the design space for the latent diffusion model, we implement a new inference method compatible with most ODE solvers.
+Applied to UniPC(Zhao et al., 2023), it can generate 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al.
+(2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively, see Fig 1.
+Figure 1: The comparison of FID performance using Free-U in different steps via ODE solvers in COCO 2014 with guidance scale $5.$ The size of the circle is proportional tothe amount of training cost.
 We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step and six-step inference exceeds most of the latest distillation models.
@@
 (2014)), and LAION datasets (Schuhmann et al.
-(2022)), and exhibit the generation result in Fig 1.
-We also use information theory to explain why our algorithm achieves a strong FID result.
-Figure 1: The comparison between our algorithm and UniPC solver in few step inference
+(2022)), and exhibit the generation results in Fig 2.
+Figure 2: The comparison between our algorithm and UniPC solver in few step inference with the prompt ”an astronaut riding a horse on Mars”.
 2 Related Work
-2.1 Diffusion ODE
+2.1 Diffusion ODE and ODE Solver
 The image-generation process of the diffusion model can be deemed as solving a special diffusion ODE.
 A typical diffusion model has two processes: forward and reverse.
-The forward process perturbs the image by adding Gaussian noise to it, which can be described by the following SDE:
-$dx=f(t)xdt+g(t)dw,$ (1)
-where the $w$ is a standard Wiener process.
-Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms.
-In this study, we use the variant preserve (VP) SDE proposed by Song et al.
-(2020b) to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm (Ho et al.
-(2020b)).
-Specifically, its choice is as follows:
-$f(t)=-\frac{1}{2}\beta(t)=\frac{d\log{\alpha_{t}}}{dt},$ (2)
-$g(t)=\sqrt{\beta(t)},$ (3)
-$\beta(t)=\beta_{min}+t(\beta_{max}-\beta_{min}).$ (4)
-Explicitly, in the diffusion process mentioned above, the variant of noise added at step $t$ is $\beta_{t}$.
-Explicilty, $\beta_{max}$ and $\beta_{min}$ can be determined arbitrarily if it can ensure $p(x_{T}|x_{0})$ is a standard Gaussian distribution.
-In the timestep $t$, the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the eq.˜5.
-$p(x_{t}|x_{0})=\mathcal{N}(\alpha(t)x_{0};\sigma^{2}(t)I).$ (5)
-Typically, to transform an image into a standard Gaussian noise, the DDPM algorithm requires $T=1000$ steps, and $\alpha_{T}$ is closed to the zero.
-Additionally, using the reparameterization algorithm, we can sample $x_{t}$ in a single step (Ho et al.
-(2020a)).
-In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE proposed by Karras et al. (2022):
-$\frac{dx}{dt}=\frac{\dot{s}(t)}{s(t)}x_{t}+s(t)^{2}\dot{\sigma}(t)\epsilon_{\theta}(x_{t},t,c),$ (6)
-where $c$ is the condition of input, like the caption that describes the image, and the $\epsilon_{\theta}$ is modeled by a neural network.
-The $\epsilon_{\theta}$ has a relation between the score function, which is described in the eq.˜7:
-$\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)score_{\theta}(x_{t},c),$
-$score_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)},$ (7)
-And the $p_{\theta}(x_{t},c)$ is the probability distribution of noisy images $x_{t}$ modeled implicitly by the neural network with parameters $\theta$.
-Moreover, the score function normally used in VP SDE is trained by eq.˜8
-[…]
-and $q_{t}(x)$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$, and $q_{0}$ means the distribution of clean images.
-In the eq.˜6, the $s(t)$ and $\sigma(t)$ is defined by eq.˜9 and eq.˜10:
-$\sigma(t)=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1},$ (9)
-$s(t)=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}},$ (10)
-where $\beta_{d}=\beta_{max}-\beta_{min}$.
-Deeming the loss of the DDPM algorithm as a score function described by the eq.˜8 can help exploit the advantage of adapting the original diffusion model to the common diffusion ODE framework proposed by Karras et al.
-(2022).
-2.2 Diffusion ODE Solver
-Sampling an image from a diffusion model is equivalent to solving the eq.˜6.
+The forward process perturbs the image by adding Gaussian noise to it.
+Hence, in the timestep $t$, the probability density of the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the Equation 1.
+$q(x_{t}|x_{0})=\mathcal{N}(\alpha_{t}x_{0};\sigma^{2}_{t}I).$ (1)
+where $\alpha_{t}$ and $\sigma_{t}$ can be choose arbitrarily if it can ensure the final $p(x_{T}|x_{0})$ close to a standard Gaussian distribution.
+In contrast, the reverse process, which corresponds to generating an image from pure Gaussian noise, can be described by a time-dependent ODE.
 Via the research from Lu et al.
-(2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the eq.˜11 :
+(2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the Equation 2 :
 $x_{t}=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\epsilon_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (2)
-where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$.
-Sometimes, to cope with the high guidance scale on the classifier free guidance algorithm (CFG) (Ho & Salimans (2022)), researchers tend to replace eq.˜11 with the eq.˜12 to get a better result (Lu et al.
-(2025)):
-$x_{t}=\frac{\sigma({t})}{\sigma(s)}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}\hat{x}_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (12)
-where $\hat{x}_{\theta}$ is a clean feature predicted indirectly by the neural network.
-By using the Taylor Expand in the $\epsilon_{\theta}$ or $x_{\theta}$ (Lu et al. (2022; 2025)), we get the following eq.˜13:
+where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$,and the $\epsilon_{\theta}$ is modeled by a neural network with ground truth proportional to the score function $\nabla_{x}\log{q(x_{t})}$, which is trained with the objective function described in Equation 3
+$\mathcal{L}_{\theta}=E_{x_{0}}[q(x_{t}|x_{0})||\frac{\epsilon_{\theta}(x_{t},t)}{\sigma_{t}}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}],$ (3)
+By using the Taylor Expand in the $\epsilon_{\theta}$ (Lu et al. (2022, 2025)), we get the following Equation 4:
 $x_{t}$ $=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\Sigma_{n=0}^{k-1}{\epsilon}^{(n)}_{\theta}(x_{t(\lambda)},\lambda)\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\frac{(\lambda-\lambda_{s})^{n}}{n!}d\lambda$
 $+\mathcal{O}(h_{i}^{k+1}),$ (4)
+where $h_{i}\propto\lambda_{t}-\lambda_{s}$
 However, calculating the neural networks’ deviation requires substantial computational resources.
-Researchers typically use the Runge-Kutta method (Lu et al.
+Researchers typically use different algorithm, like the Runge-Kutta method (Lu et al.
 (2025); Zhao et al.
-(2023)) to approximate the deviation.
-Explicitly, the DPM++ 2s algorithm approximate $x_{t}$ via the following formula:
-$x_{t}$ $=\frac{\sigma({t})}{\sigma(s)}x_{s}-\alpha_{t}(e^{-h_{t}}-1)\hat{x}_{\theta}(x_{s},s)$
-$+\frac{1}{2r_{t}}*\alpha_{t}(e^{-h_{t}}-1)*(\hat{x}_{\theta}(\mu,m)-\hat{x}_{\theta}(x_{s},s)),$ (14)
-in which $r_{t}$ can be selected arbitarily between 0 and 1, and $h_{i}$, $m$ and $\mu_{i}$ is given below:
-$h_{t}$ $=\lambda_{t}-\lambda_{s},$ (15)
-$r_{t}$ $=\frac{\lambda_{m}-\lambda_{s}}{h_{i}},$ (16)
-$\mu_{t}$ $=\frac{\sigma_{m}}{\sigma_{s}}x_{s}-\alpha_{m}(e^{-r_{t}h_{t}}-1)\hat{x}_{\theta}(x_{s},s),$ (17)
-To get the first-order deviation of the $\epsilon_{\theta}$, the number of function evaluations (NFE) in DPM++ 2s is 2.
-While DPM++ 2m utilizes the latest cached output in the previous inference to calculate the approximation of the first-order deviation, and hence only requires 1 NFE, which is a typical multistep algorithm in ODE solving.
-Simultaneously, the UniPC and DPM V3 solver cached all of the previous output to correct and predict a high-order deviation.
-The eq.˜6 describes a continuous curved trajectory, and the ODE solver approximates it by discretizing the path with a series of lines.
+(2023)), to approximate the deviation.
+2.2 Time Schedule
+The original diffusion ODE describes a continuous curved trajectory, and the Equation 2 approximates it by discretizing the path with a series of lines.
 Research on reverse ODE solvers aims to precisely estimate the next step, thereby minimizing truncation error from pool discretization in the few-step inference.
 On the other hand, the discrete method proposed by Karras et al.
-(2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ to enhance the sample quality.
+(2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ at each step $t$ to enhance the sample quality.
 Via the analysis of the Karras’ research, the $h_{t}$ should be large at the beginning of the inference, and then decrease monotonically.
+Since truncation error is bound by the Equation 5
+[…]
+where $\mathcal{T}_{i}||\leq Ch_{i}$.
+The $E$ depends on the discrete step $N$, the start time step $t_{0}$, the ending timestep $t_{N}$, and the existing constant $C$.
+The meaning of this upper bound is limited by the part that contributes the highest $\mathcal{T}_{i}$.
+And Karras et al.
+(2022) reveals that the truncation error of the diffusion model increases monotonously when the $h_{i}$ in each inference step is equal.
+Based on this observation, Karras et al.
+(2022) proposes a following discrete method:
+$\sigma_{i}=(\sigma_{min}^{\frac{1}{p}}+\frac{i}{N-1}(\sigma_{min}^{\frac{1}{p}}-\sigma_{max}^{\frac{1}{p}}))^{p}.$ (6)
+Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget.
+Based on that observation, Li & Zhang (2025) proposes another discrete method:
+$t_{i}=(t_{min}^{\frac{1}{p}}+\frac{i}{N-1}(t_{min}^{\frac{1}{p}}-t_{max}^{\frac{1}{p}}))^{p}.$ (7)
+When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method.
+Hence, it can address the problem with the original Karras schedule mentioned above.
 2.3 Diffusion Distillation
@@
 (2024); Luo et al.
-; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as an extensive neural network.
+; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as a huge neural network.
 The distillation algorithm, which typically trains a small neural network, reduces its inference steps.
 However, training a neural network with high-quality content, fast inference, and no additional parameters is difficult.
-Thus, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance.
-This approach may pose a potential risk in the future.
-AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules.
-However, when a human draws an image via their craft, they will sometimes be affected by the image they have previously seen.
-If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such a generation will not be deemed plagiarism.
-Compared to this, generative models trained on Internet images are often criticized, with people claiming that if those models generate content similar to their dataset, it should be considered plagiarism.
-In such a situation, diversity is a metrics that affect people’s impression of generative content.
+Actually, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance.
+We do further analysis and experiments to enhance the augmentation.
 2.4 Latent Diffusion Model
 Another method for reducing the computational requirements of inference is to encode images into a low-dimensional latent space.
-Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image.
-One experiment from Rombach et al.
-(2022) shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism (Vaswani et al.
-(2017)).
-In contrast, its perceptual information, which provides the image with sufficient real details, is strongly tied to local regions and can be modeled by a convolutional neural network (CNN) (e.g.
-Li et al.
-(2021)).
-Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al.
+Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 image.
+To achieve this, it firstly encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al.
 (2018)) to let the following inference ignore the perceptual detail.
@@
 3 Approach
-3.1 Algorithm Based on Truncation Error Analysis
-To design a training-free, few-step inference algorithm, we should first consider the meaning of the original diffusion models’ output, the score function.
-And then compare it with the distillation model’s output.
-By rewriting the eq.˜8 as follows:
-$\mathcal{L}_{\theta}=\frac{E_{x_{0}}[q(x_{t}|x_{0})||s_{\theta}-\nabla_{x}\log q(x_{t}|x_{0})||^{2}]}{E_{x_{0}}[q(x_{t}|x_{0})]},$ (18)
-We find out that the score function is pointing to the average of all possible images that can cause this noisy image.
+3.1 The effect of the skip-connection modification
+Modifying the skip connection via Free-U(Si et al., 2024), a training-free U-Net decorator, can reduce the low-frequency component of the noisy skip connection and enhance the backbone feature, thereby augmenting the denoising capability of the diffusion model while preserving image details(Si et al., 2024).
+We find that the skip connection modification removes the blur from the generative image, implying a change in the neural network’s original output from a standard score function to a vector pointing to a special image, which allows the model’s behaviour to be similar to that of the diffusion disillation model.
+However, experiments in Figure 3 show that Free-U will also change the discrete method of the diffusion ODE, and this effect cannot be ignored in the few-step inference that has a poor discrete situation.
+Explicitly, the output of the neural network in Equation 3 is pointing to the average of all possible images that can cause this noisy image.
 In an extreme few-step inference case, like one or two NFE inference, the final output will be blurred, since in that step, the score function only points to the centroid of images.
 Meanwhile, the distillation algorithm, like rectify flow proposed by Liu (2022), will tend to modify the output and let it point to a special sample.
-Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, one can presume that the score function $\nabla_{x}\log q(x_{t})$ can be approximated by the conditional score function $\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage, since the possible images are very limited.
-In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
-Since the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, one of the natural method is to enhance the denoise capability of U-Net, which can be achieved by modifying the skip connection and the backbone feature of the U-Net.
-Previous study (Si et al.
-(2024)) adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability.
-Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the low-frequency part of the noisy backbone feature map.
-The decorator only applies to the first two skip connections.
-The scaler that affects the backbone feature is $b_{1}$ and $b_{2}$, while the scaler that affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
-However, we don’t know which step could be changed simply from the score function to the conditional score function, and one of the vanilla ways is to apply free-U at all time steps.
+Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, the Free-U that preserves the detail of the output image while augmenting the denoise capability can partly achieve it in the least noisy stage, since the possible images are very limited.
+The previous research from Ma et al.
+(2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers.
+However, we don’t know which step could be changed simply from the score function to the conditional score function pointing to special a sample, and one of the vanilla ways is to apply free-U at all time steps, as the original paper claims.
 This method has been seen as helpful to slightly modify the score function and make it more similar to the distillation’s output.
-The previous research from Ma et al.
-(2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers, since the skip connection itself will abate the result in the block below that layer, while the few-step inference tends to excute a huge transportation from a standard Gaussian distribution to a complex probability distribution, and hence require the skip connection reduce its’ effect that abate the block below it.
-However, via the mathematical analysis of the diffusion ODE and the experimental result.
+But, via the mathematical analysis of the diffusion ODE and the experimental result in Figure 4.
 This vanilla method in the extreme few-step inference of high-resolution images, such as 6 NFE, would not work properly.
 Explicitly, consider that the normal ResNet (He et al. (2016)) can be viewed as an ODE (Chen et al. (2018)) by using the following formula:
-$h_{t+1}=h_{t}+f_{\theta}(h_{t},t),$ (19)
-$\frac{dh}{dt}=F_{\theta}(h_{t},t),$ (20)
-where $h_{t}$ is the feature output in layer $t$.
-The U-Net, as a black box ODE solver, can be regarded as another ODE.
-Thus, modifying the skip connection and scale backbone feature will further increase the norm of the output, increasing the moving distance of the diffusion reverse ODE.
-In the situation of the few-step inference, it will be a serious problems, since the weakness of the poor discretization method will be amplified by this behaviour.
-The related mathematic details will be mentioned below.
-We also visualize the decoded features along the trajectories of different samplers in fig.˜2.
-Figure 2: The comparison between our custom DPM++1s sampler and the original DPM++1s sampler.
-The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler.
-The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
-To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by applying the same sampler DPM++ 1s sampler in fig.˜3.
-The standard time schedule using Free-U generates a trajectory similar to that of Karras et al.
-(2022) time schedule without Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference, which determines the image layout composition.
+$l_{t+1}=l_{t}+f_{\theta}(l_{t},t),$ (8)
+$\frac{dl}{dt}=F_{\theta}(l_{t},t),$ (9)
+where $l_{t}$ is the feature output in layer $t$.
+The U-Net, as a black box ODE solver using skip connections, can be regarded as another ODE.
+And modifying the skip connection and scale backbone feature in the neural ODE will further increase the $h_{i}$ in Equation 4.
+To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by utilizing the same sampler DPM++ 1s sampler in Figure 3.
+The uniform time schedule using Free-U generates a feature in the first two steps that is similar to the Karras et al.
+(2022) time schedule without Free-U, while totally different to the normal uniform sampler without using Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference.
+And if we use karras or a similar time schedule, it will have an unwilling effect, since it’s large enough at the beginning step.
 Figure 3: The comparison between different discrete methods when using the same DPM++ 1s Sampler in a few-step inference.
 The first line exhibits the trajectory of the normal Karras discrete method.
-The second one exhibits the trajectory of the Karras discrete method + Free-U.
+The second one exhibits the trajectory of the Karras discrete method + Free-U, which has an unwilling result.
 The third exhibits the normal discrete method with a normal U-Net, and the final one exhibits the normal discrete method with Free-U.
 And the final line is similar to the first line.
-Another method is to parameterize the start step of using the U-Net decorator in an existing solver.
+To prevent this problem, another method is to parameterize the start step of using the U-Net decorator in an existing solver.
 However, experimental results show that this Vallina method doesn’t work either.
-We visualize the 6-step inference result combining the free-U and UniPC solver in fig.˜4.
+We visualize the 6-step inference result combining the free-U and UniPC solver in Figure 4.
 Figure 4: The result that starts using free-U in the different steps of the UniPC solver.
-To address black-box ODE-solving problems, we should identify the characteristics of the diffusion ODE.
-The basic logic of the ODE solver is to discretize a continuous curve trajectory into a series of lines.
-And to achieve the best performance, the ODE solver should determine the length of each discrete line for different ODEs, such as the Fehlberg method (see e.g.
-Atkinson et al.
-(2009)) that adjusts the discrete length based on the error per unit stepsize.
-For the diffusion ODE, when the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as fatal as the non-noisy steps.
-These insights have been proved by Karras et al.
-(2022) via an experiment that discretizes the inference process with a uniform gap between each timestep.
-And it obtains an array of steps $\{t_{i}\}$.
-Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$.
-The experimental results show that the truncation error increases monotonically as the noise variance add in the input feature at the time step $t_{i}$ decreases.
-[…]
-Based on this observation, Karras et al. (2022) proposes a following discrete method:
-$\sigma_{i}=(\sigma_{min}^{\frac{1}{p}}+\frac{i}{N-1}(\sigma_{min}^{\frac{1}{p}}-\sigma_{max}^{\frac{1}{p}}))^{p}.$ (21)
-eq.˜21 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part.
-Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget.
-Considering the U-Net has 860 million parameters, and the variance of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical.
-Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
-Based on that observation, Li & Zhang (2025) proposes another discrete method:
-$t_{i}=(t_{min}^{\frac{1}{p}}+\frac{i}{N-1}(t_{min}^{\frac{1}{p}}-t_{max}^{\frac{1}{p}}))^{p}.$ (22)
-When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method.
-In such a situation, $t_{i}-t_{i-1}$ will increase in the noisier part.
-The eq.˜4 shows that $t_{i}$ is proportional to the variance of noise added in $x_{t}$, which can be deemed as the speed of movement at timestep $t_{i}$.
-The increase of $t_{i}-t_{i-1}$ in the noisy part aligns with the rise in speed, while the rise of $\sigma_{i}-\sigma_{i-1}$ aligns with the increment of the moving distance, hence it can cope with the problem of the original Karras schedule mentioned above.
-However, both of those discrete methods have been commonly used in all kinds of images.
-Our research shows that, even with the same time schedule and UniPC solver, perceptual quality in the few-step inference remains substantially different across image resolution and image type.
-Hence, the design space of hyperparameters for different ODE solvers has not yet been totally exploited.
-We find a discrete method, which changes the eq.˜23, a formula with a special design reason mentioned later, by adding one more step to get clean denoised images.
-And use it to generate images at different resolutions and of different types in a few-step inference.
-We find that the result in fig.˜5 has a noticeable gap in perceptual quality.
-Figure 5: The comparison between the different resolution images and different type images.
-The first line contains 512 x 512 images generated by Stable Diffusion 1.5, the second line contains 1024 x 1024 images generated by Dreamshaper XL.
-The last lines contain high-resolution ACGN images generated by Illustrij Evo.
-Based on the experimental results, designing a customized training-free algorithm for different image types requires exploring the design space of the ODE solver for the latent diffusion model.
-Firstly, the decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget.
+To address black-box ODE-solving problems, we should provide a new time schedule.
+3.2 The time schedule leveraging $\beta$-VAE
+The decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget.
 Although the latent diffusion model theoretically does not require using the KL term to train an autoencoder, this optimization trick is very prevalent in the training of the diffusion model.
-The latest representation autoencoders (RAE) also add noise during the training to augment the decoder Zheng et al.
-(2025).
+The latest representation autoencoders (RAE) also add noise during the training to augment the decoder (Zheng et al., 2025).
 The experimental results show that the decoder’s capacity to handle noisy features varies with image resolution.
-In the fig.˜6, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem.
-And the third line output feature with less noise, and has a better results.
+In the Figure 5, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem.
+And the third line output feature has less noise and produces better results.
 Figure 5: The comparison between the different resolution images generated by the same time schedule and ODE solver utilizing the same $\beta$-VAE.
@@
 $t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}},$ (10)
-where $\sigma_{stop}$ is defined via the following formula:
+where $\sigma_{stop}$ is defined via Equation 11, and $t(\sigma)$ is a function projecting variant $\sigma$ to timestep $t$, which use the discrete version describe in Equation 1 whose range is between $[1,1000]$:
 $\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{M}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (11)
-which could be nonzero to save the one-step inference budget.
-By applying this special time schedule, the problems of using Free-U while preventing an increase in moving distance could be addressed, since we could only apply Free-U at timesteps that move a short distance.
+and $\sigma$ could be nonzero to save the one-step inference budget by utilizing $\beta$-VAE.
+Meanwhile, the problems of using Free-U while preventing an increase in moving distance could be addressed by applying this special time schedule, since we could only apply Free-U at timesteps that move a short distance.
 By leveraging the robustness of the $\beta$-VAE, the total moving distance has also been constrained.
-The visual difference between our discrete method and the method proposed by the Karras’ paper is in fig.˜7
+The visual difference between our discrete method applying on DPM++ 2m, and the method proposed by the Karras’ paper is in Figure 6
 Figure 6: The comparison between our time schedule and the time schedule proposed by Karras
-Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) during inference.
-Our algorithm can be described by the algorithm˜1
+Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) when generating high-resolution images.
+Our algorithm can be described by the Algorithm 1
 Data: number of inference step $N$, initiated noise $\epsilon$, stop step in Karras schedule $t_{stop}$, the total step $M$ of the original discrete method, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$, a bool value afsafs to determine whether using afs, and an analytical first step solver without using neural network $F$
 Result: final result image
-1 $\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$;
-2
-[…]
-4if afsafs then
+$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{t_{stop}}{N+t_{stop}+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}}$
+$ts=\{t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}};0\leq i<N\}$
+if afsafs then
 2 $ts.insert((\frac{t_{0}+t_{1}}{2}),at:1)$
 3 end if
-7
 […]
-9while $t$ in tsts do
 5 if $t$ equal to $t_{aug}$ then $freeU(F_{\theta})$  ;
-11   ;
 6 if afsafs and $t$ equal to 999999 then $x=F(x)$ ;
 7 else $x=F_{\theta}(x)$ ;
-14
-15 end while
-Algorithm 1 custom Karras scheduler in the common framework of Diffusion ODE solving
+8
+9end while
+Algorithm 1 F-scheduler in the common framework of Diffusion ODE solving
 3.3 Information Theory Analysis of the Different Algorithms
-The experimental results in the section˜4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance.
-Given that all images are generated from the same prompts, and the distillation algorithm changes the score function from pointing to the centroid of all possible images to a special image, we assume the distillation model limits the diversity of images’ low-frequency components, which is caused by modification of the score function at the start stage of the inference.
-Explicitly, some images with totally different semantics could have a similar, fuzzy structure with the same color block in a similar position.
-And the limitation on the possible fuzzy structure of images may also cause diversity problems.
-And we use information theory to explain why applying Free-U only in the final stage of the inference helps address this problem.
-To introduce the relationship between information theory and machine learning, we should first review the historical confusion about machine learning’s generalization capability.
-A common point of confusion in deep learning at an early stage is how neural networks with so many parameters can still generalize (Breiman (2018)).
-The standard statistical model uses a limited number of parameters to concisely describe the phenomenon and capture the essence of the observed data, whereas deep learning is fundamentally different.
-For this problem, the study theory (Harman & Kulkarni (2012)) claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions.
-Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{N}$.
-According to information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can access is $2^{H(D)}$ (Yeung (2012)).
-For example, in a high-dimensional standard Gaussian distribution, a pure black image, although with the highest possible density, is not in the typical set.
-Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
-Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$.
-The information bottleneck theory proposed by Tishby et al.
-(2000) shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ projected to the $Z$.
-Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses by limiting the possible middle feature (Tishby et al.
-(2000)).
-As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) (Goodfellow et al.
-(2020)), which face mode collapse (Creswell et al.
-(2018)) that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint Jeon et al.
-(2021).
-One study (Li & Zhang (2025)) shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information.
-The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE).
-When $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bound of the mutual information constraint will be abated.
-Meanwhile, the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$ will increase.
-From this perspective, the diffusion model can be deemed as a hierarchy VAE (Luo (2022)), while progressive distillation proposed by Salimans & Ho (2022) increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$.
-Where $X_{\theta}$ is a noisy feature output by solving ODE.
-Hence, the PSNR of the middle output feature also increases.
-However, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will also affect the mutual information, even if the output does not have an increased PSNR.
+The experimental results in the Section 4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance.
+To explain why our algorithm can outperform the distillation model, we use information theory to provide a tentative analysis of the diffusion model.
+Bertrand et al.
+(2025) shows that using the ground truth score function in the place having the highest loss, explicitly, 20% of inference, to replace the studied score function will harm the diversity of the diffusion model.
+They don’t provide a theoretical analysis of this phenomenon, nor do they provide a further estimate of the distillation model that modifies the score function.
+In this section, the information bottleneck theory Theorem 3.1 reveals that mutual information affects the model’s generalization.
+We find out when the score function been changed from the centroid of all possible images that can generate the current noisy image to a special image, the mutual information between the middle feature and the output increases, as described in Proposition 3.2.
+By preventing the mutual information from affecting, Theorem 3.1 and Proposition 3.3 show that the current average output of the model would not shift and deviate from the ground truth centroid, while the mode collapse in GAN, shown in Figure 7, is different from that.
+Jeon et al.
+(2021) also shows that constraining the information bottleneck in GAN can enhance the FID.
+Eplicilty, for an arbitrary image generative model, we assume that there is a ground truth noise $x$ and image $y$ pair that is possible to be implicitly studied by the neural network, and $y$ sample from $\mathcal{Y}$ contains images with enough diversity.
+Then, according to learning theory, if we want to prevent a model from rote, we should constrain the number of functions this model can express, to ensure it can get access to the best hypothesis in training.
+Suppose we have a training data $s_{n}=\{y_{i}\}_{i=1}^{n}$ that has at least $1-\sigma$ probability drawn from the same distribution of $\mathcal{Y}$, a neural network $f^{s}$ trained in $s$ via SGD in $i$ iteration, a batch size $m$ and a batch validating data $s_{m}=\{\hat{y}_{i}\}_{i=1}^{m}$, a middle feature $Z_{l}^{s}$ output from the layer $l$ of the neural network $f^{s}$.
+Then, for the output of the $f^{s}$, we consider a output $f^{s}(\hat{x}_{1}),f^{s}(\hat{x}_{2}),...,f^{s}(\hat{x}_{m})]$ from a network on $i$-th iteration traning which the $m$ noise-image pair $\hat{x},\hat{y}$ is denote by the ground truth matching.
+We consider the finite hypothesis set $\mathcal{H}$ that consists of all possible outputs in the $i$-th iteration of SGD training, without considering the order of outputs within the same batch.
+Meanwhile, in SGD, the training process is random, and with a given noise-feature pair $x,z_{l}^{s}$ in training, we assume $p(Z_{l}^{s}|Y)$, who denote the probability density of $Z_{l}^{s}$ in the $i$ iteration of SGD training, will use $z_{l}^{s}$ as the means.
+And the mutual information $I(Y;Z_{l}^{s})$ is given by $p(Z_{l}^{s},Y)$.
+Finally, seperate the neural network into two parts: decoder $g^{s}$ and encoder $\phi^{s}$, and $f^{s}=g^{s}\circ\phi^{s}$, where we use $\circ$ to denote the composition of functions.
+First, the following theorem investigates the information bottleneck (IB) of the latent diffusion model, whose proof is listed in the Appendix A.
+Theorem 3.1.
+For the L2 loss function $\mathcal{L}$, and $\Delta(s)=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]-\frac{1}{m}\Sigma_{i=1}^{m}\mathcal{L}(f^{s}(x_{i}),y_{i})$, we have:
+$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};s_{n})}+log\frac{2}{\sigma}}{2o*A_{m}^{m}}},$ (12)
+where $o$ is a fixed constant for every hypothesis space, and $A_{m}^{m}$ denotes the number of arrangements of $m$ distinct objects.
+Unlike (Kawaguchi et al., 2023) that applies IB to a stochastic neural network, we also show that, even without the stochastic part in the neural network, the SGD itself will give the mutual information between the middle feature and the input, and it still constrains the possible hypotheses.
+Add a stochastic process to the neural network, or use the loss with high variance, just a special case of our analysis.
+Then, it is intuitive that changing the direction of the score function to allow it to point to the final result, rather than the centroid of the possible image, will affect the mutual information, then affect the Equation 19, and cause the diversity problem.
 We provide further analysis regarding this.
-The diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following eq.˜25:
-$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))$
-$\leq H(X(0))-E_{Z_{\theta}(t)}[p(Z_{\theta}(0)|Z_{\theta}(t))$
-$*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))],$ (25)
-in which $H(X(0))$ is the entropy of the real $X(0)$ sample from the dataset.
-In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$, enhancing the log likelihood of the $Y$.
-And in the distillation algorithm, eq.˜25 shows that if we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase.
-Most distillation methods that do not increase the PSNR of the output will still face this problem.
-We evaluate the Precision and Recall of Distribution (PRD) (Sajjadi et al.
-(2018)) of flash diffusion and our sampler to further validate this claim.
+Explicitly, the diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following Proposition 3.2:
+Proposition 3.2.
+$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))\leq H(Z(0))-E_{Z_{\theta}(t)}[-p(Z_{\theta}(0)|Z_{\theta}(t))*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))]$
+in which $H(Z(0))$ is the entropy of the real $Z(0)$ sample from the dataset.
+If we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase.
+In the following proposition, we decompose the $E_{X,Y}[\mathcal{L}]=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]$ into two parts to explain mode collapse.
+And $\Delta(s)$ been constrain by IB above is the uppper bound of the $E_{X,Y}[\mathcal{L}]$
+Proposition 3.3.
+$E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]=E_{X,Y}[\mathcal{L}]-E_{X,Y}[(\hat{f}(X)-y)^{2}]$, where $\hat{f}(x)=E_{s_{n}}[f^{s}(x)]$
+Meanwhile, as the Proposition 3.2 explain, the constrain of mutual information also reduce the log likelihood related to the second term.
+And, in GAN, the mode collapse will enlarge the $E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]$, like the Figure 7.
+Hence Proposition 3.3 can prevent the diversity problem.
+The proof of Proposition 3.3 is postponed to Appendix B.
+Figure 7: Illustration of Mode Collapse.
+(Left) The target distribution consists of a mixture of 8 Gaussians arranged in a ring.
+(Right) A simulated example of mode collapse that contains only a subset of the available modes.
 4 Experiments
 4.1 Ablation Study
-We conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampling from the COCO 2017 dataset.
+To demonstrate the necessity of applying Free-U at different steps, we conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampled from the COCO 2017 dataset.
 Explicitly, for the 8-step inference, we employ $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ as the discrete setting.
@@
 Don’t apply Free-U.
-The experimental result is in the table˜1.
-The $w$ in the table˜1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
+The experimental result is in the Table 1.
+The $w$ in the Table 1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
 Meanwhile, for the 8-step UniPC solver, we use a 2nd-order multistep algorithm, whereas the 6-step solver is 1st-order solver.
-Solver FID CLIP-Score
-custom UniPC(6) apply Free-U on the every steps $w=5.5$ 28.88 30.96
-custom UniPC(8) apply Free-U on the every steps $w=5.5$ 33.70 30.50
-custom UniPC(6) don’t apply Free-U $w=5.5$ 27.13 31.49
-custom UniPC(8) don’t apply Free-U $w=5.5$ 25.47 31.39
-custom UniPC(6) start apply Free-U at 4th step $w=5.5$ 25.20 31.42
-custom UniPC(8) start apply Free-U at 6th step $w=5.5$ 24.38 31.38
-custom UniPC(6) apply Free-U on the every steps $w=7.5$ 39.62 30.33
-custom UniPC(8) apply Free-U on the every steps $w=7.5$ 54.26 29.63
-custom UniPC(6) don’t apply Free-U $w=7.5$ 25.34 31.53
-custom UniPC(8) don’t apply Free-U $w=7.5$ 24.61 31.22
-custom UniPC(6) start apply Free-U at 4th step $w=7.5$ 25.29 31.44
-custom UniPC(8) start apply Free-U at 6th step $w=7.5$ 24.58 31.21
 Table 1: Results of applying Free-U in different steps of inference.
+Solver steps use Free-U apply stage FID CLIP-Score
+UniPC 6 Yes Every Steps 28.88 30.96
+UniPC 8 Yes Every Steps 33.70 30.5
+UniPC 6 No Never 27.13 31.49
+UniPC 8 No Never 25.47 31.39
+UniPC 6 Yes 4th step 25.20 31.42
+UniPC 8 Yes 6th step 24.38 31.38
 4.2 Performance
@@
 We employ $p_{1}=7$, $p_{2}=1.2$, $stop=8$, $M=12$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images via DPM++ 2m.
-Additionally, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images.
-And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$.
-Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$.
-And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$.
-Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
-For the eight-step inference that uses DPM++ 2m to generate 512x512-resolution images, we avoid using free-U.
-And for the customized DPM++ 2m solver that generates 1024x1024-resolution images, the $t_{aug}$ apply free-U is 8.
-For all UniPC solver customized to generate 512x512 resolution images, $t_{aug}=4$.
-And $t_{aug}=6$, $t_{aug}=4$ for 6-step UniPC solver and 8-step UniPC solver customeized for 1024 x 1024-resolution images, respectively.
-And we only use AFS on the UniPC solver customized for the 1024 x 1024-resolution image.
-We don’t guarantee we provide the best hyperparameters; we just ensure we provide hyperparameters that can sample images with FID performance comparable to that of the inference result from the 20-step DPM++ solver and the SOTA distillation algorithm.
-We report the results of the inference that generate 512 x 512-resolution images in the COCO 2014, COCO 2017, and LAION dataset in table˜2, table˜3, and table˜4.
-Moreover, the result of the 1024 x 1024 images is given in table˜5, table˜6, and table˜7.
+And its $t_{aug}$ is -1 and 8 for 512x512 and 1024x1024 resolutions, respectively.
+Additionally, $t_{aug}=6$, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images.
+And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $t_{aug}=4$, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$.
+Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $t_{aug}=4$, $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$.
+And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $t_{aug}=4$, $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$.
+Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $t_{aug}=4$, $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
+We report the results of the inference that generate images via 10000 captions in the COCO 2014, 5000 captions in COCO 2017, and 10000 captions in the LAION dataset in Table 2, Table 5, Table 3, Table 6, Table 4, and Table 7.
 Explicitly, the UniPC solver is a 2nd-order multistep solver for eight-step inference and a 1st-order multistep solver for five and six-step inference.
 The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively.
-Furthermore, all the synthesized images in table˜2 and table˜5 employs the random 10k COCO caption as a guidance prompt.
-In table˜3 and table˜6, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance.
-Furthermore, we use 10k captions in LAION for table˜4 and table˜7.
 We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2 (Yin et al.
@@
 (2024)) are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison.
-The $w$ in the table˜2 and table˜5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
+The $w$ in the Table 2 and Table 5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model.
 The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
-Solver FID CLIP-Score
-(Our) custom DPM++1s(5) $w=7.5$ 19.18 30.92
-(Our) custom DPM++1s(6) $w=7.5$ 18.54 31.16
-(Our) custom DPM++2m(8) $w=7.5$ 15.92 31.32
-(Our) custom DPM++1s(5) $w=5.5$ 17.54 31.0
-(Our) custom DPM++1s(6) $w=5.5$ 17.04 31.16
-(Our) custom DPM++2m(8) $w=5.5$ 15.7 31.07
-(Our) custom UniPC(5) $w=7.5$ 19.35 30.63
-(Our) custom UniPC(6) $w=7.5$ 16.41 30.97
-(Our) custom UniPC(8) $w=7.5$ 16.00 31.15
-(Our) custom UniPC(5) $w=5.5$ 15.19 30.93
-(Our) custom UniPC(6) $w=5.5$ 14.72 31.08
-(Our) custom UniPC(8) $w=5.5$ 15.26 31.12
-Original DPM++2m(20) $w=7.5$ 17.3 30.97
-Original DPM++2m(20) $w=5.5$ 18.68 30.97
-Flash Diffusion(5) 17.34 30.45
-Flash Diffusion(6) 17.94 30.59
-Flash Diffusion(8) 19.03 30.37
-AMED Pulgin(8) 19.07 31.15
-SDXL turbo(4) 23.24 31.66
 Table 2: Results of 512 x 512 image generation. COCO 2014
-Solver FID CLIP-Score
-(Our) custom DPM++1s(5) $w=5.5$ 24.22 31.06
-(Our) custom DPM++1s(6) $w=5.5$ 23.24 31.12
-(Our) custom DPM++2m(8) $w=5.5$ 22.40 31.23
-(Our) custom DPM++1s(5) $w=7.5$ 25.58 30.93
-(Our) custom DPM++1s(6) $w=7.5$ 24.71 31.13
-(Our) custom DPM++2m(8) $w=7.5$ 22.35 31.35
-(Our) custom UniPC(5) $w=7.5$ 22.70 30.99
-(Our) custom UniPC(6) $w=7.5$ 24.02 31.04
-(Our) custom UniPC(8) $w=7.5$ 24.55 30.99
-(Our) custom UniPC(5) $w=5.5$ 21.31 30.92
-(Our) custom UniPC(6) $w=5.5$ 20.82 31.04
-(Our) custom UniPC(8) $w=5.5$ 21.71 31.07
-Original DPM++2m(20) $w=7.5$ 25.97 31.21
-Original DPM++2m(20) $w=5.5$ 23.75 31.10
-Flash Diffusion(5) 23.56 30.64
-Flash Diffusion(6) 24.16 30.60
-Flash Diffusion(8) 25.35 30.55
-AMED Pulgin(8) $w=7.5$ 25.50 31.17
-SDXL turbo(4) 30.92 31.66
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 15.92 31.32
+DPM++ 2m 8 5.5 No Never 15.7 31.07
+UniPC 5 7.5 Yes 4-th step 19.35 30.63
+UniPC 6 7.5 Yes 4-th step 16.41 30.97
+UniPC 8 7.5 Yes 4-th step 16.00 31.15
+UniPC 5 5.5 Yes 4-th step 15.19 30.93
+UniPC 6 5.5 Yes 4-th step 14.72 31.08
+UniPC 8 5.5 Yes 4-th step 15.26 31.12
+DPM++ 2m 20 7.5 No Never 17.3 30.97
+DPM++ 2m 20 5.5 No Never 18.68 30.97
+Flash Diffusion 5 None No Never 17.34 30.45
+Flash Diffusion 6 None No Never 17.94 30.59
+Flash Diffusion 8 None No Never 19.03 30.37
+AMED Pulgin 8 7.5 No Never 19.07 31.15
+SDXL turbo 4 None No Never 23.24 31.66
 Table 3: Results of 512 x 512 image generation. COCO 2017
-Solver FID CLIP Score
-(Our) custom DPM++1s(5) $w=5.5$ 19.61 31.81
-(Our) custom DPM++1s(6) $w=5.5$ 18.45 32.10
-(Our) custom DPM++2m(8) $w=5.5$ 17.74 32.42
-(Our) custom DPM++1s(5) $w=7.5$ 19.88 31.45
-(Our) custom DPM++1s(6) $w=7.5$ 18.60 31.97
-(Our) custom DPM++2m(8) $w=7.5$ 17.52 32.41
-(Our) custom UniPC(5) $w=7.5$ 20.43 30.87
-(Our) custom UniPC(6) $w=7.5$ 17.56 31.63
-(Our) custom UniPC(8) $w=7.5$ 17.38 32.08
-(Our) custom UniPC(5) $w=5.5$ 17.44 31.73
-(Our) custom UniPC(6) $w=5.5$ 16.84 32.08
-(Our) custom UniPC(8) $w=5.5$ 17.32 32.26
-Original DPM++2m(20) $w=7.5$ 18.30 32.67
-Original DPM++2m(20) $w=5.5$ 17.33 32.47
-Flash Diffusion(5) 15.40 31.60
-Flash Diffusion(6) 15.80 31.63
-Flash Diffusion(8) 16.87 31.57
-AMED Pulgin(8) $w=7.5$ 18.06 32.44
-SDXL turbo(4) 22.25 33.02
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 22.40 31.23
+DPM++ 2m 8 5.5 No Never 22.35 31.35
+UniPC 5 7.5 Yes 4-th step 22.70 30.99
+UniPC 6 7.5 Yes 4-th step 24.02 31.04
+UniPC 8 7.5 Yes 4-th step 24.55 30.99
+UniPC 5 5.5 Yes 4-th step 21.31 30.92
+UniPC 6 5.5 Yes 4-th step 20.82 31.04
+UniPC 8 5.5 Yes 4-th step 21.71 31.07
+DPM++ 2m 20 7.5 No Never 25.97 31.21
+DPM++ 2m 20 5.5 No Never 23.75 31.10
+Flash Diffusion 5 None No Never 23.56 30.64
+Flash Diffusion 6 None No Never 24.16 30.60
+Flash Diffusion 8 None No Never 25.35 30.55
+AMED Pulgin 8 7.5 No Never 25.50 31.17
+SDXL turbo 4 None No Never 30.92 31.66
 Table 4: Results of 512 x 512 image generation. Laion
-Solver FID CLIP Score
-custom DPM++2m (8) $w=7.5$ 17.84 31.62
-custom DPM++2m (8) $w=5.5$ 19.62 31.58
-custom UniPC (8) $w=7.5$ 18.18 31.23
-custom UniPC (8) $w=5.5$ 17.87 31.40
-custom UniPC (6) $w=7.5$ 18.80 31.42
-custom UniPC (6) $w=5.5$ 19.07 31.45
-Flash DiffusionXL (6) 22.23 31.00
-Flash DiffusionXL (8) 23.25 30.89
-SDXL-Lightning (8) 21.07 31.11
-DMD2 (4) 19.64 31.64
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 17.74 32.42
+DPM++ 2m 8 5.5 No Never 17.52 32.41
+UniPC 5 7.5 Yes 4-th step 20.43 30.87
+UniPC 6 7.5 Yes 4-th step 17.56 31.63
+UniPC 8 7.5 Yes 4-th step 17.38 32.08
+UniPC 5 5.5 Yes 4-th step 17.44 31.73
+UniPC 6 5.5 Yes 4-th step 16.84 32.08
+UniPC 8 5.5 Yes 4-th step 17.32 32.26
+DPM++ 2m 20 7.5 No Never 18.30 32.67
+DPM++ 2m 20 5.5 No Never 17.33 32.47
+Flash Diffusion 5 None No Never 15.40 31.60
+Flash Diffusion 6 None No Never 15.80 31.63
+Flash Diffusion 8 None No Never 16.87 31.57
+AMED Pulgin 8 7.5 No Never 18.06 32.44
+SDXL turbo 4 None No Never 22.25 33.02
 Table 5: Results of 1024 x 1024 image generation. COCO 2014
-Solver FID CLIP Score
-custom DPM++2m (8) $w=7.5$ 24.42 31.71
-custom DPM++2m (8) $w=5.5$ 26.04 31.56
-custom UniPC (8) $w=7.5$ 24.58 31.21
-custom UniPC (8) $w=5.5$ 24.38 31.38
-custom UniPC (6) $w=7.5$ 25.29 31.44
-custom UniPC (6) $w=5.5$ 25.20 31.42
-Flash DiffusionXL (6) 28.74 30.96
-Flash DiffusionXL (8) 29.63 30.88
-SDXL-Lightning (8) 27.89 31.08
-DMD2 (4) 26.19 31.65
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 17.84 31.62
+DPM++ 2m 8 5.5 No Never 19.62 31.58
+UniPC 8 7.5 Yes 4-th step 18.18 31.23
+UniPC 8 5.5 Yes 4-th step 17.87 31.40
+UniPC 6 7.5 Yes 4-th step 18.80 31.42
+UniPC 6 5.5 Yes 4-th step 19.07 31.45
+Flash Diffusion 6 None No Never 22.23 31.00
+Flash Diffusion 8 None No Never 23.25 30.89
+SDXL-Lightning 8 None No Never 21.07 31.11
+DMD2 4 None No Never 19.64 31.64
 Table 6: Results of 1024 x 1024 image generation. COCO 2017
-Solver FID CLIP Score
-custom DPM++2m (8) $w=7.5$ 19.25 32.22
-custom DPM++2m (8) $w=5.5$ 20.82 32.07
-custom UniPC (8) $w=7.5$ 21.59 31.73
-custom UniPC (8) $w=5.5$ 20.41 32.06
-custom UniPC (6) $w=7.5$ 22.51 31.82
-custom UniPC (6) $w=5.5$ 22.33 31.87
-Flash DiffusionXL (6) 20.88 31.64
-Flash DiffusionXL (8) 21.30 31.61
-SDXL-Lightning (8) 20.45 32.50
-DMD2 (4) 16.51 33.28
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 24.42 31.71
+DPM++ 2m 8 5.5 No Never 26.04 31.56
+UniPC 8 7.5 Yes 4-th step 24.58 31.21
+UniPC 8 5.5 Yes 4-th step 24.38 31.38
+UniPC 6 7.5 Yes 4-th step 25.29 31.44
+UniPC 6 5.5 Yes 4-th step 25.20 31.42
+Flash Diffusion 6 None No Never 28.74 30.96
+Flash Diffusion 8 None No Never 29.63 30.88
+SDXL-Lightning 8 None No Never 27.89 31.08
+DMD2 4 None No Never 26.19 31.65
 Table 7: Results of 1024 x 1024 image generation.Laion
+Solver steps scale use Free-U apply stage FID CLIP-Score
+DPM++ 2m 8 7.5 No Never 19.25 32.22
+DPM++ 2m 8 5.5 No Never 20.82 32.07
+UniPC 8 7.5 Yes 4-th step 21.59 31.73
+UniPC 8 5.5 Yes 4-th step 20.41 32.06
+UniPC 6 7.5 Yes 4-th step 22.51 31.82
+UniPC 6 5.5 Yes 4-th step 22.33 31.87
+Flash Diffusion 6 None No Never 20.88 31.64
+Flash Diffusion 8 None No Never 21.30 31.61
+SDXL-Lightning 8 None No Never 20.45 32.50
+DMD2 4 None No Never 16.51 33.28
 We also calculate the PRD in a 6-step inference by sampling 10k captions from COCO 2014 and getting the corresponding image in the COCO dataset as a reference.
-PRD was first used to evaluate GAN.
-Considering two distributions $Q$ and $P$, the first one is the distribution of the generative images, and the second one is the distribution of the original images, the precision shows how much of $Q$ can be generated by $P$, and the recall exhibits how much $P$ can be generated by $Q$.
-For example, if $P$ is a subset of MNIST that only contains numbers ranging from 1 to 4, adding 5 and 6 to the generative distribution reduces the precision, while lacking modes like 3 and 4 will reduce the recall rate.
-Our experiment shows that the recall rate of the flash diffusion in the 6-step is less than that of us.
+Our experiment shows that the generative diversity of the flash diffusion in the 6-step is less than that of us.
 The guidance scale of our sampler is 5.
@@
 In this study, we propose a plug-in method to enhance the sampling algorithm.
-By fully harnessing the potential capacity of the latent diffusion model—such as modifying the skip-connection at the right time, using the property of the decoder, which is a $\beta$-VAE trained with small KL-constraints, ignoring a small amount of noise in the final feature to reschedule the discretizing method, and choosing the proper ODE solver via truncation error analysis—we can sample a high-quality image with limited inference budget and avoid the need for additional training.
+By fully harnessing the potential capacity of the latent diffusion model—such as modifying the skip-connection at the right time, using the property of the decoder, ignoring a small amount of noise in the final feature to reschedule the discretizing method, and choosing the proper ODE solver via truncation error analysis—we can sample a high-quality image with limited inference budget and avoid the need for additional training.
 5.1 Limitation and the future work
-Firstly, a diffusion model is a robust generative model (Chen et al.
-(2024)) that can prevent adversarial attacks.
-Most attacks that hack the latent diffusion model attack its VAE decoder.
-Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that largely depends on the property of the VAE.
-Moreover, in future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
-Last, the hyperparameters of the ACGN model need further exploration.
+In future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
+Appendix A Proof of Theorem 3.1
+Firstly, consider the well-known probably approximately correct (PAC)(Hafez-Kolahi et al., 2020) deduction from Hoeffding’s inequality, we have:
+$\Delta(s)=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]-\frac{1}{m}\Sigma_{i=1}^{m}\mathcal{L}(f^{s}(x_{i}),y_{i})\leq\sqrt{\frac{\log{|\mathcal{H}|+\log{\frac{2}{\sigma}}}}{2o}},$ (13)
+with probability approaching 1, where $o$ is a given constant for each hypothesis space
+And input $X^{(m)}=\{x_{1},x_{2},...,x_{m}\}$, in which sample $x_{i}$ samples i.i.d from $\mathcal{X}$, is a sample from the $\mathcal{X}^{(m)}$.
+The $|\mathcal{H}|\leq 2^{|\mathcal{X}^{(m)}|}$.
+Then, for the typical set $\mathcal{A}_{\epsilon}^{(m)}$ of the $\mathcal{X}^{(m)}$, that the probability of $X^{(m)}$ in $\mathcal{A}_{\epsilon}^{(m)}$ is bigger than $1-\epsilon$, and its cardinality $|\mathcal{A}_{\epsilon}^{(m)}|\leq 2^{m(H(x)+\epsilon)}$, then, when the batch size itself is large and the $\epsilon$ is small enough, we have:
+$\Delta(s)\leq\sqrt{\frac{2^{mH(X)}+\log{\frac{2}{\sigma}}}{2o}}$ (14)
+Where $H(X)$ is the entropy of the $X$.
+Then, seperate the neural network into two parts: decoder $g^{s}$ and encoder $\phi^{s}$, and $f^{s}=g^{s}\circ\phi^{s}$, where we use $\circ$ to denote the composition of functions.
+And we use $h^{s}_{i}$ to denote the $i$-th layer of the $n$-layer neural network.
+Explicilty, $\phi^{s}_{l}=h^{s}_{l}\circ h^{s}_{l-1}\circ\dots\circ h^{s}_{1}$ and $g^{s}_{l}=h^{s}_{n}\circ h^{s}_{n-1}\circ\dots\circ h^{s}_{l+1}$.
+Then, the cardinality of the hypothesis space of the decoder $g^{s}_{l}$ is constrained by the number of input features $z_{l}^{s}$.
+And for the hypothesis space $\mathcal{H}_{g}$ with input $Z_{l}^{s}$:
+$\Delta(s)\leq\sqrt{\frac{2^{mH(Z_{l}^{s})}+\log{\frac{2}{\sigma}}}{2o}}$ (15)
+And then we consider the conditional typical set given by the $p(Z_{l}^{s}|Y)$.
+When the $Y$ is given, the cardinality of the typical set of $\{{Z_{l}^{s}}^{(1)},{Z_{l}^{s}}^{(1)},\dots,{Z_{l}^{s}}^{(m)}\}$ is $2^{mH(Z_{l}^{s}|Y)}$.
+Hence for each decoder $g^{s}_{l}$ with input $Z_{l}^{s}$ that having a mutual information constrain $I(Z_{l}^{s};Y)$, the total possible middle feature set is equal to:
+$|\mathcal{H}_{g}|=\frac{2^{H(Z_{l}^{s})}}{2^{H(Z_{l}^{s}|Y)}}=2^{I(Z_{l}^{s};Y)}$ (16)
+Then, for each decoder $g^{s}_{l}$, we have
+$\Delta(s_{g})\leq\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+log\frac{2}{\sigma}}{2o}}$ (17)
+Meanwhile, for each encoder $\phi_{l}^{s}$, we use $I(\phi_{l}^{s};S)$ to constrain the overfitting.
+Explicitly, the cardinality of the hypothesis space $\mathcal{H_{\phi}}$ of the encoder is $2^{I(\phi_{l}^{s};S)}$ (Kawaguchi et al., 2023).
+And then, we mulitiply the cardinality of two hypothesis spaces $\mathcal{H}_{g}$ and $\mathcal{H}_{\phi}$, to get the hypothesis space of $f_{l}^{s}$ whose the decoder is constrained with mutual information, and the encoder is in the $\mathcal{H_{\phi}}$, we have:
+$\Delta(s_{l})\leq\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o}}$ (18)
+And the $\Delta(s)$, when each encoder is constrained by a mutual information $I(\phi_{l}^{s};S)$ to prevent overfiting, is:
+$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o}}$ (19)
+Then, if we ignore the order of output in a batch of validation data, we have
+$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o*A_{m}^{m}}}$ (20)
+and A is the arrangement number, which means the mutual information gives a bound on the generalization error. $\Box$
+Appendix B Proof of Proposition 3.3
+For the generalization error $E_{X,Y}[{\mathcal{L}(f^{s}(X),Y)}]$, we have
+$\mathcal{L}(f^{s}(x),y)$ $=(f^{s}(x)-y)^{2}$ (21)
+$=(f^{s}(x)-E[y|x]+E[y|x]-y)^{2}$ (22)
+$=(f^{s}(x)-E[y|x])^{2}+(E[y|x]-y)^{2}+2(f^{s}(x)-E[y|x])(E[y|x]-y).$ (23)
+Recall that
+$E_{X,Y}[\mathcal{L}]$ $=\int_{x}\int_{y}\mathcal{L}(f^{s}(x),y)p(x,y)dxdy.$ (24)
+By (24), we compute first and third term on the right hand side of (23):
+$\int_{x}\int_{y}(f^{s}(x)-E[y|x])^{2}p(x,y)dxdy$ $=\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx,$ (25)
+$\int_{x}\int_{y}2(f^{s}(x)-E[y|x])(E[y|x]-y)dxdy$ $=0.$ (26)
+Therefore, we have
+$E_{X,Y}[\mathcal{L}]$ $=\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx+\int_{x}\int_{y}(E[y|x]-y)^{2}p(x,y)dxdy,$ (28)
+where
+$\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx$ $=E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]+E_{X,Y}[(\hat{f}(X)-E[y|x])^{2}]$ (29)
+where $E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]$ denote the distance of the output from a particular training neural network $f^{s}$ and the average output $\hat{f}$. Note that
+$\int_{x}\int_{y}(E[y|x]-y)^{2}p(x,y)dxdy+E_{X,Y}[(\hat{f}(X)-E[y|x])^{2}]=E_{X,Y}[(\hat{f}(X)-y)^{2}].$ (30)
+Combining (28), (29) and (30), we derive the final result. $\Box$